### v1

# Phantom Crossing and Oscillating Dark Energy with $F(R)$ Gravity

Abstract
In this work, we shall consider how a dynamical oscillating and phantom crossing dark energy era can be realised in the context of $F(R)$ gravity. We approach the topic from a theoretical standpoint, considering all the conditions that may lead to a consistent phantom crossing behaviour and separately how the $F(R)$ gravity context may realise an oscillating dark energy era. Apart from our qualitative considerations, we study quantitatively two $F(R)$ gravity dark energy models, which are viable cosmologically and also exhibit simultaneously phantom crossing behaviour and oscillating dark energy. We consider these models by numerically solving the field equations using appropriate statefinder parameters engineered for dark energy studies. As we show, $F(R)$ provides a natural extension of Einstein’s general relativity, which can naturally realise a transition from a phantom era to a quintessential era, a feature supported by recent observational data, without resorting to phantom scalar fields to realise the phantom evolution.
preprint: KEK-TH-2734, KEK-Cosmo-0383
I Introduction
The observation of the late-time acceleration of the Universe, known as dark energy (DE) era is one of the most unexpected features of our world. In the context of simple general relativity, the DE era can be generated by quintessence scalar fields, however, the latest observations of the Dark Energy Spectroscopic Instrument (DESI) DESI:2024mwx ; DESI:2025zgx indicate that the DE was in the phantom regime in the near past. This feature can be described in the context of general relativity only by using phantom scalar fields, which is a rather unattractive feature. Also and more importantly, the DESI data indicate a transition from a phantom to a quintessential regime for DE. This is questionable if it can be realised in the context of general relativity. There are many ways to produce results compatible with the DESI data, for example Odintsov:2024woi ; Dai:2020rfo ; He:2020zns ; Nakai:2020oit ; DiValentino:2020naf ; Agrawal:2019dlm ; Ye:2020btb ; Vagnozzi:2021tjv ; Desmond:2019ygn ; Hogas:2023pjz ; OColgain:2018czj ; Vagnozzi:2019ezj ; Krishnan:2020obg ; Colgain:2019joh ; Vagnozzi:2021gjh ; Lee:2022cyh ; Krishnan:2021dyb ; Ye:2021iwa ; Ye:2022afu ; Verde:2019ivm ; Menci:2024rbq ; Adil:2023ara ; Reeves:2022aoi ; Ferlito:2022mok ; Vagnozzi:2021quy ; DiValentino:2020evt ; Sabogal:2024yha ; Giare:2024smz ; DiValentino:2025sru ; Odintsov:2025kyw ; vanderWesthuizen:2025iam ; Paliathanasis:2025dcr ; Odintsov:2025jfq ; Kessler:2025kju ; Nojiri:2025jhh ; Ozulker:2025ehg and among these frameworks, $F(R)$ gravity Nojiri:2017ncd ; Capozziello:2011et ; Nojiri:2006ri ; Nojiri:2010wj ; Nojiri:2003ft ; Capozziello:2005ku ; Hwang:2001pu ; Song:2006ej ; Faulkner:2006ub ; Olmo:2006eh ; Sawicki:2007tf ; Faraoni:2007yn ; Carloni:2007yv ; Nojiri:2007as ; Deruelle:2007pt ; Appleby:2008tv ; Dunsby:2010wg ; Oikonomou:2025qub enjoys an elevated role. There are recent studies which discuss the ability of $F(R)$ gravity to be compatible with the DESI data Odintsov:2025jfq . Another more fascinating behaviour of DE based on the DESI data was demonstrated in Ref. Kessler:2025kju , which indicates that the DE might not be only dynamical and evolving from phantom to quintessence, but it also shows strong indications of an oscillating equation of state (EoS). In this work, we aim to investigate whether a phantom involving and oscillating DE may be realised by $F(R)$ gravity in general. We approach the problem theoretically, finding the conditions which may allow such a phenomenological behaviour to be produced, and also we present several phenomenological models from a theoretical point of view. In addition, we study numerically the DE era of some explicit $F(R)$ gravity models which can realise a viable Planck compatible Planck:2018vyg and in some cases DESI compatible late-time behaviour. We also demonstrate that these models produce an oscillating DE era and also allow for phantom to quintessence transitions.
This paper is organised as follows: In Section II, we discuss the inverse-phantom processing in $F(R)$ gravity from a theoretical viewpoint. The analysis is continued in Section III, where we discuss the apparent phantom crossing. In Section IV, the theoretical study of several models that exhibit phantom crossing is considered. The oscillating Hubble rate is considered in Section V, while in Section VI, the effects of a scalaron as dark matter with curvature-dependent mass on the DE era are considered. In Section VII, we discuss the features of a realistic model, while in Section VIII, we numerically study two explicit examples of oscillating $F(R)$ gravity DE with phantom crossing. Finally, the conclusions follow at the end of the article.
II Inverse-Phantom Crossing in $F(R)$ gravity
Consider the vacuum $F(R)$ gravity action,
$\mathcal{S}=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}F(R)\,,$ (1)
and a flat Friedmann-Lemaître-Robertson-Walker (FLRW) metric with the following line element,
$\centering ds^{2}=-dt^{2}+a(t)\sum_{i=1}^{3}dx_{i}^{2}\,.\@add@centering$ (2)
Let $w$ be the DE equation of state (EoS) parameter, that is, the ratio of the pressure of the DE fluid with respect to the energy density of the DE fluid. If the recent DESI observations DESI:2024mwx ; DESI:2025zgx are correct, there might be a transition from $w<-1$ to $w>-1$ in the DE EoS.
In this section, we discuss if the (a little bit initial) condition determining if the standard phantom crossing (from $w>-1$ to $w<-1$) or the inverse phantom crossing (from $w<-1$ to $w>-1$) can be realised theoretically in the context of $F(R)$ gravity. Note that for future phantom crossing in $F(R)$ gravity, one can consult Bamba:2008hq . We show that the (inverse) phantom crossing in the $F(R)$ sector, which is described by the effective pressure and the effective energy density, can occur for almost any $F(R)$ gravity by adjusting the (initial) conditions.
The conservation law for a FLRW spacetime is,
$0=\dot{\rho}+3H\left(\rho+p\right)\,,$ (3)
indicates that when the phantom crossing occurs, since we have $\rho=-p$, then $\dot{\rho}$ must vanish. Due to the fact that in the phantom Universe, $\rho<-p$, $\dot{\rho}$ and in the non-phantom Universe, $\rho>p$, there is a crossing from the non-phantom Universe to the phantom one if $\ddot{\rho}>0$ and there is an inverse crossing if $\ddot{\rho}<0$.
Let the $F(R)$ gravity be of the form,
$F(R)=R+f(R)\,.$ (4)
The effective energy density $\rho_{f(R)}$ and the effective pressure $p_{f(R)}$ coming from the $f(R)$ part are given by,
$\rho_{f(R)}=$ $\,\frac{1}{\kappa^{2}}\left(-\frac{1}{2}f(R)+3\left(H^{2}+\dot{H}\right)f^{\prime}(R)-18\left(4H^{2}\dot{H}+H\ddot{H}\right)f^{\prime\prime}(R)\right)\,,$ (5)
$p_{f(R)}=$ $\,\frac{1}{\kappa^{2}}\left(\frac{1}{2}f(R)-\left(3H^{2}+\dot{H}\right)f^{\prime}(R)+6\left(8H^{2}\dot{H}+4{\dot{H}}^{2}+6H\ddot{H}+\dddot{H}\right)f^{\prime\prime}(R)\right.$
$\,\left.+36\left(4H\dot{H}+\ddot{H}\right)^{2}f^{\prime\prime\prime}(R)\right)\,.$ (6)
We consider the phantom crossing, where $\dot{\rho}_{f(R)}=\rho_{f(R)}+p_{f(R)}=0$, in the $f(R)$ sector by using (5) and (II). The Friedmann equation is given by,
$\frac{3}{\kappa^{2}}H^{2}=\rho_{f(R)}+\rho_{\mathrm{matter}}\,,$ (7)
where we also considered the contribution of perfect matter fluids. We now assume that the matter energy density $\rho_{\mathrm{matter}}$ is given by the dust, that is, the baryonic matter and the cold dark matter. Then we may assume,
$\rho_{\mathrm{matter}}=\rho_{\mathrm{matter(0)}}a^{-3}=\rho_{\mathrm{matter(0)}}\mathrm{e}^{-3N}\,.$ (8)
Here $\rho_{\mathrm{matter(0)}}$ is a constant and $a$ is the scale factor, which is given in terms of the $e$-foldings $N$ as $a=\mathrm{e}^{N}$.
When the phantom crossing in the $f(R)$ sector happens at the cosmic time instance $t=t_{\mathrm{ph}}$, we assume $N$ can be expanded as,
$N=\sum_{n=0}\frac{H_{n}}{\left(n+1\right)!}\left(t-t_{\mathrm{ph}}\right)^{n+1}\,,$ (9)
which gives the following Hubble rate $H$,
$H=\sum_{n=0}\frac{H_{n}}{n!}\left(t-t_{\mathrm{ph}}\right)^{n}\,.$ (10)
Then we find that the scalar curvature $R$ behaves as,
$R=$  $\,12H^{2}+6\dot{H}=12{H_{0}}^{2}+6H_{1}+\left(24H_{0}H_{1}+6H_{2}\right)\left(t-t_{\mathrm{ph}}\right)$
$\,+\left(12H_{0}H_{2}+12{H_{1}}^{2}+3H_{3}\right)\left(t-t_{\mathrm{ph}}\right)^{2}+\mathcal{O}\left(\left(t-t_{\mathrm{ph}}\right)^{3}\right)\,.$ (11)
We also assume that,
$\rho_{f(R)}=\rho_{{f(R)}\,(0)}+\rho_{{f(R)}\,(2)}\left(t-t_{\mathrm{ph}}\right)^{2}+\mathcal{O}\left(\left(t-t_{\mathrm{ph}}\right)^{3}\right)\,.$ (12)
Then if $\rho_{{f(R)}\,(2)}>0$, a crossing from the non-phantom $\rho_{f(R)}>p_{f(R)}$ to the phantom $\rho_{f(R)}<p_{f(R)}$ occurs and if $\rho_{{f(R)}\,(2)}<0$, an inverse crossing occurs.
The first Friedmann equation (7) indicates that,
$\frac{3}{\kappa^{2}}{H_{0}}^{2}=$ $\,\rho_{{f(R)}\,(0)}+\rho_{\mathrm{matter(0)}}\,,$ (13)
$\frac{6}{\kappa^{2}}H_{0}H_{1}=$ $\,-3\rho_{\mathrm{matter(0)}}H_{0}\,,$ (14)
$\frac{3}{\kappa^{2}}\left(H_{0}H_{2}+2{H_{1}}^{2}\right)=$ $\,\rho_{{f(R)}\,(2)}+\rho_{\mathrm{matter(0)}}\left(\frac{9}{2}{H_{0}}^{2}-\frac{3}{2}H_{1}\right)\,.$ (15)
On the other hand, Eq. (5) tells,
$\rho_{f(R)\,(0)}=$ $\,\frac{1}{\kappa^{2}}\Bigl{(}-\frac{1}{2}f\left(12{H_{0}}^{2}+6H_{1}\right)+3\left({H_{0}}^{2}+H_{1}\right)f^{\prime}\left(12{H_{0}}^{2}+6H_{1}\right)$
$\,-18\left(4{H_{0}}^{2}H_{1}+H_{0}H_{2}\right)f^{\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)\Bigr{)}\,,$ (16)
$0=$ $\,\frac{1}{\kappa^{2}}\left\{-6H_{0}H_{1}f^{\prime}\left(12{H_{0}}^{2}+6H_{1}\right)\right.$
$\,+18\left(4{H_{0}}^{3}H_{1}-3{H_{0}}^{2}H_{2}-4H_{0}{H_{1}}^{2}-H_{0}H_{3}\right)f^{\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)$
$\,\left.-108H_{0}\left(4H_{0}H_{1}+H_{2}\right)^{2}f^{\prime\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)\right\}\,,$ (17)
$\rho_{f(R)\,(2)}=$ $\,\frac{1}{\kappa^{2}}\left\{-9H_{0}H_{2}f^{\prime}\left(12{H_{0}}^{2}+6H_{1}\right)\right.$
$\,+\left(-36{H_{0}}^{2}H_{2}+108H_{0}H_{1}H_{2}-144{H_{1}}^{3}+9{H_{2}}~{}2\right)f^{\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)$
$\,+\left(24H_{0}H_{1}+6H_{2}\right)\left(36{H_{0}}^{3}H_{1}-63{H_{0}}^{2}H_{2}-108H_{0}{H_{1}}^{2}-18H_{0}H_{3}-9H_{1}H_{2}\right)$
×$\,\times f^{\prime\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)$
$\,\left.-324H_{0}\left(4H_{0}H_{1}+H_{2}\right)^{3}f^{\prime\prime\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)\right\}\,.$ (18)
In the following, we consider the solution of the above equations and their physical meanings. Eq. (14) tells,
$H_{1}=-\frac{\kappa^{2}}{2}\rho_{\mathrm{matter(0)}}\,,$ (19)
therefore, $H_{1}$ is always negative. This expression can also be obtained by combining the first and second Friedmann equations,
$-\frac{2}{\kappa^{2}}\dot{H}=\rho_{f(R)}+p_{f(R)}+\rho_{\mathrm{matter}}+p_{\mathrm{matter}}\,,$ (20)
and noting $\rho_{f(R)}+p_{f(R)}=0$ when the phantom crossing occurs for $f(R)$ sector and $p_{\mathrm{matter}}=0$ for cold matter and baryonic matter.
By combining (13), (II), and (19), we find the following expression of $H_{2}$,
$H_{2}=$ $\,-4H_{0}H_{1}-\frac{f\left(12{H_{0}}^{2}+6H_{1}\right)}{36H_{0}f^{\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)}+\frac{\left({H_{0}}^{2}+H_{1}\right)f^{\prime}\left(12{H_{0}}^{2}+6H_{1}\right)}{6H_{0}f^{\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)}$
$\,-\frac{3{H_{0}}^{2}+H_{1}}{18H_{0}f^{\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)}\,.$ (21)
Eq. (15) with (19) and (II) indicate that,
$\rho_{{f(R)}\,(2)}=$ $\,\frac{3}{\kappa^{2}}\left(H_{0}H_{2}+2{H_{1}}^{2}\right)+\frac{2}{\kappa^{2}}H_{1}\left(\frac{9}{2}{H_{0}}^{2}-\frac{3}{2}H_{1}\right)$
$=$ $\,\frac{3}{\kappa^{2}}\left(-{H_{0}}^{2}H_{1}+{H_{1}}^{2}\right)-\frac{f\left(12{H_{0}}^{2}+6H_{1}\right)}{12\kappa^{2}f^{\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)}$
$\,+\frac{\left({H_{0}}^{2}+H_{1}\right)f^{\prime}\left(12{H_{0}}^{2}+6H_{1}\right)}{2\kappa^{2}f^{\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)}-\frac{3{H_{0}}^{2}+H_{1}}{6\kappa^{2}f^{\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)}\,.$ (22)
Therefore, it depends on the detailed form of $f(R)$ if a crossing from the non-phantom to the phantom occurs, that is, $\rho_{{f(R)}\,(2)}>0$, or an inverse crossing, that is, $\rho_{{f(R)}\,(2)}<0$, occurs. This could indicate that the phantom crossing or the inverse phantom crossing occurs in a wide class of $F(R)$ gravity models, depending on the initial condition determining $H_{0}$. Furthermore, by combining (II) with (II), we can find the expression of $H_{3}$, although we do not write it explicitly because we do not use the expression of $H_{3}$ later.
As an example, we may consider $R^{2}$ gravity,
$f(R)=\alpha R^{2}\,.$ (23)
Here $\alpha$ is a constant. As is well-known, the $F(R)$ gravity includes a scalar mode. In the model (23), the square of the mass is given by $\frac{1}{\alpha}$. Therefore, in order to avoid the tachyon, we require $\alpha$ to be positive.
For the model (23), Eq. (II) has the following form,
$\rho_{{f(R)}\,(2)}=$ $\,\frac{1}{2\kappa^{2}}\left(9{H_{1}}^{2}-\frac{3{H_{0}}^{2}+H_{1}}{6\alpha}\right)\,.$ (24)
For the standard phantom crossing to occur, the condition is given by
${H_{0}}^{2}<18\alpha{H_{1}}^{2}-\frac{1}{3}H_{1}=\frac{9\alpha\kappa^{4}}{2}{\rho_{\mathrm{matter(0)}}}^{2}+\frac{\kappa^{2}}{6}\rho_{\mathrm{matter(0)}}\,,$ (25)
were we used Eq. (19). On the other hand, the condition for the inverse phantom crossing is given by,
${H_{0}}^{2}>18\alpha{H_{1}}^{2}-\frac{1}{3}H_{1}=\frac{9\alpha\kappa^{4}}{2}{\rho_{\mathrm{matter(0)}}}^{2}+\frac{\kappa^{2}}{6}\rho_{\mathrm{matter(0)}}\,.$ (26)
Then the crossing depends on the initial condition to determine $H_{0}$ and $\rho_{\mathrm{matter(0)}}$.
III Apparent Phantom Crossing
There is an idea that instead of considering the transition from $w<-1$ to $w>-1$ of DE, the modification of the dark matter sector may solve the problem in the DESI observations Khoury:2025txd . If dark matter density decreases more slowly than $a^{-3}$, as usually predicted from energy conservation, it looks like there could be a phantom crossing in the DE sector because we are considering the total energy density, and we assume the usual $a^{-3}$ behaviour of the dark matter for the analysis of the DESI observations.
If the dark matter is non-relativistic, as in the cold dark matter, the energy density of the dark matter is given by the product $\rho_{\mathrm{DM}}=m_{\mathrm{DM}}n_{\mathrm{DM}}$ of the mass of the dark matter particle, $m_{\mathrm{DM}}$, and the number density of the dark matter particle, $n_{\mathrm{DM}}$. If there is neither creation nor annihilation of the dark matter particles, the number density $n_{\mathrm{DM}}$ is proportional to the inverse of the volume in the expanding Universe, $n_{\mathrm{DM}}\propto V^{-1}\propto a^{-3}$. If the mass $m_{\mathrm{DM}}$ is time dependent $m_{\mathrm{DM}}(t)$ or $m_{\mathrm{DM}}$ depends on the scalar curvature $R$, the behaviour of the dark matter energy density $\rho_{\mathrm{DM}}$ is different from $a^{-3}$. If the mass depends on the scalar curvature, the energy density also includes the derivative terms because the curvature includes the derivative, which makes the situation a little bit complicated.
In order to make the physical problem at hand well-defined, we assume that the dark matter is a Dirac spinor, whose action is given by
[…]
$\mathcal{H}=\bar{\psi}\left(-i\gamma^{i}D_{i}+m(R)\right)\psi\,,$ (28)
in the canonical formulation. The energy density is the quantum or statistical average of the Hamiltonian density, as follows,
[…]
The first term corresponds to the kinetic energy and the second term to the rest mass. As long as we consider non-relativistic particles, the first term can be neglected by comparing the second term. We should also note that $\bar{\psi}\psi\right>$ is nothing but the number density of dark matter $n_{\mathrm{DM}}$. Then we obtain,
$\rho_{\mathrm{DM}}=m(R)n_{\mathrm{DM}}\,,$ (30)
as expected. Hence, if we choose $m(R)$ to be a decreasing function of $R$, at late times where $R$ decreases, $m(R)$ becomes larger and $\rho_{\mathrm{DM}}$ decreases slower than $a^{-3}$.
The $R$-dependence of the mass, makes the situation complicated when we consider the variation with respect to the metric or the vierbein field. The variation of the action (27) with respect to the metric or the vierbein field gives the following energy-momentum tensor of the Dirac field,
[…]
$\,+2\left(-R_{\mu\nu}-g_{\mu\nu}\Box+\nabla_{\mu}\nabla_{\nu}\right)\left(m^{\prime}(R)\bar{\psi}\psi\right)\,.$ (31)
On the other hand, the Dirac equations given by the action (27) with respect to the Dirac spinors, $\psi$ and $\bar{\psi}$ have the following forms,
[…]
We assume that $\bar{\psi}\gamma_{i}D_{j}\psi\right>$ and $\left(D_{i}\bar{\psi}\right)\gamma_{j}\psi\right>$ are negligible for the non-relativistic dark matter. Then we find the energy density $\tilde{\rho}_{\mathrm{DM}}$ and the pressure $\tilde{p}_{\mathrm{DM}}$ as follows,
$\tilde{\rho}_{\mathrm{DM}}=$ $\,m(R)\left<\bar{\psi}\psi\right>+\left\{3\left(\dot{H}+H^{2}\right)+3H\frac{d}{dt}\right\}\left(m^{\prime}(R)\left<\bar{\psi}\psi\right>\right)$
$=$ $\,m(R)n_{\mathrm{DM}}+\left\{3\left(\dot{H}+H^{2}\right)+3H\frac{d}{dt}\right\}\left(m^{\prime}(R)n_{\mathrm{DM}}\right)\,,$
$\tilde{p}_{\mathrm{DM}}=$$\,\left\{-\left(\dot{H}+3H^{2}\right)+\frac{d^{2}}{dt^{2}}+2H\frac{d}{dt}\right\}\left(m^{\prime}(R)\left<\bar{\psi}\psi\right>\right)$
$=$ $\,\left\{-\left(\dot{H}+3H^{2}\right)+\frac{d^{2}}{dt^{2}}+2H\frac{d}{dt}\right\}\left(m^{\prime}(R)n_{\mathrm{DM}}\right)\,.$ (33)
The expression of $\tilde{\rho}_{\mathrm{DM}}$ in (III) is different from $\rho_{\mathrm{DM}}$ in (30) by the second term.
By combining $\tilde{\rho}_{\mathrm{DM}}$ and $\tilde{p}_{\mathrm{DM}}$ with $\rho_{f(R)}$ and the effective pressure $p_{f(R)}$ coming from $f(R)$ term in (5) and (II), we find the total effective energy density $\rho_{\mathrm{total}}\equiv\tilde{\rho}_{\mathrm{DM}}+\rho_{f(R)}$ and the total pressure $p_{\mathrm{total}}\equiv\tilde{p}_{\mathrm{DM}}+p_{f(R)}$, which satisfy the standard conservation law, $0=\dot{\rho}_{\mathrm{total}}+3H\left(\rho_{\mathrm{total}}+p_{\mathrm{total}}\right)$. We may subtract the part which behaves as $a^{-3}$ from the effective energy density $\rho_{\mathrm{eff\,DE}}$ of the DE,
$\rho_{\mathrm{eff\,DE}}\equiv\rho_{\mathrm{total}}-\rho_{0}a^{-3}\,.$ (34)
Here $\rho_{0}$ is a constant corresponding to the energy density of the dark matter in the present Universe if we choose $a=1$ in the present Universe. We may also identify the effective pressure $p_{\mathrm{eff\,DE}}$ of DE with the total pressure $p_{\mathrm{total}}$, $p_{\mathrm{eff\,DE}}=p_{\mathrm{total}}$. Then the effective energy and the effective pressure satisfy the conservation law, $0=\dot{\rho}_{\mathrm{eff\,DE}}+3H\left(\rho_{\mathrm{eff\,DE}}+p_{\mathrm{eff\,DE}}\right)$. Then even if the effective energy density $\rho_{f(R)}$ and the pressure $p_{f(R)}$ of the $f(R)$ sector do not show the phantom crossing, the effective energy density $\rho_{\mathrm{eff\,DE}}$ and the effective pressure $p_{\mathrm{eff\,DE}}$ of DE may appear to cause the phantom crossing.
IV Several Models that Exhibit Phantom Crossing: A Theoretical Account
We now consider some specific examples that exhibit DE phantom crossing. We may consider the model Odintsov:2024woi ,
$F(R)=R-2\Lambda\left(1-\mathrm{e}^{-\beta\left(\frac{R}{2\Lambda}\right)^{\alpha}}\right)\,,$ (35)
where $\Lambda$ is a constant with dimension $[m]^{2}$ and $\beta$ is a dimensionless constant. When $R$ is large, $\beta\left(\frac{R}{2\Lambda}\right)^{\alpha}\gg 1$ if $\alpha>0$, and $\Lambda$ plays the role of the cosmological constant. This model is also proposed to solve the inverse phantom crossing. We choose $\frac{\Lambda}{\beta^{\frac{1}{\alpha}}}$ small enough compared with the present value of the scalar curvature. This makes $\Lambda$ give the DE in the present Universe. In order to also describe the inflation in the early Universe, we may add the term $\gamma R^{2}$ in the $F(R)$ of (35) with a constant $\gamma$.
Because $f(R)=-2\Lambda\left(1-\mathrm{e}^{-\beta\left(\frac{R}{2\Lambda}\right)^{\alpha}}\right)$, we find
$f^{\prime}(R)=$ $\,-\alpha\beta\left(\frac{R}{2\Lambda}\right)^{\alpha-1}\mathrm{e}^{-\beta\left(\frac{R}{2\Lambda}\right)^{\alpha}}\,,$
$f^{\prime}(R)=$ $\,-\frac{\alpha\beta}{2\Lambda}\left\{\left(\alpha-1\right)\left(\frac{R}{2\Lambda}\right)^{\alpha-2}-\alpha\beta\left(\frac{R}{2\Lambda}\right)^{2\alpha-2}\right\}\mathrm{e}^{-\beta\left(\frac{R}{2\Lambda}\right)^{\alpha}}\,.$ (36)
Then $\rho_{{f(R)}\,(2)}$ in (II) has the following form,
$\rho_{{f(R)}\,(2)}=$ $\,\frac{3}{\kappa^{2}}\left(-{H_{0}}^{2}H_{1}+{H_{1}}^{2}\right)-\frac{\Lambda^{2}\left(1-\mathrm{e}^{-\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha}}\right)\mathrm{e}^{\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha}}}{3\alpha\beta\kappa^{2}\left\{\left(\alpha-1\right)\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-2}-\alpha\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{2\alpha-2}\right\}}$
$\,+\frac{\Lambda\left({H_{0}}^{2}+H_{1}\right)\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-1}}{\kappa^{2}\left\{\left(\alpha-1\right)\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-2}-\alpha\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{2\alpha-2}\right\}}$
$\,+\frac{\Lambda\left(3{H_{0}}^{2}+H_{1}\right)\mathrm{e}^{\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha}}}{3\alpha\beta\kappa^{2}\left\{\left(\alpha-1\right)\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-2}-\alpha\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{2\alpha-2}\right\}}\,.$ (37)
Then, depending on the choice of the values of $H_{0}$ and $H_{1}$ or dark matter energy density, the phantom crossing or the inverse phantom crossing occurs. Especially, we choose the simple case $\alpha=1$. Then Eq. (IV) reduces to,
$\rho_{{f(R)}\,(2)}=$ $\,\frac{3}{\kappa^{2}}\left(-{H_{0}}^{2}H_{1}+{H_{1}}^{2}\right)-\frac{\Lambda^{2}}{3\beta^{2}\kappa^{2}}\left(1-\mathrm{e}^{-\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)}\right)\mathrm{e}^{\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)}$
$\,-\frac{\Lambda}{\beta\kappa^{2}}\left({H_{0}}^{2}+H_{1}\right)-\frac{\Lambda}{3\beta^{2}\kappa^{2}}\left(3{H_{0}}^{2}+H_{1}\right)\mathrm{e}^{\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)}\,.$ (38)
Then for example, when $\Lambda$ is small enough, $\rho_{{f(R)}\,(2)}\to\frac{3}{\kappa^{2}}\left(-{H_{0}}^{2}H_{1}+{H_{1}}^{2}\right)>0$. We should note $H_{1}$ is negative due to (19). Then, in the case that $\Lambda$ is small enough, the standard phantom crossing could occur. On the other hand, when $\Lambda$ is large enough, we find $\rho_{{f(R)}\,(2)}\to-\frac{\Lambda^{2}}{3\beta^{2}\kappa^{2}}\left(1-\mathrm{e}^{-\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)}\right)<0$, and as long as $-H_{1}<2{H_{0}}^{2}$, the inverse phantom crossing could occur.
Let us consider more the formulation in Section III. By using (5) and (III), the Friedmann equation in (7) has the following form,
$0=$ $\,\frac{1}{2\kappa^{2}}\left(R+f(R)\right)-m(R)n_{\mathrm{DM}}$
+$\,+\left\{3\left(\dot{H}+H^{2}\right)+3H\frac{d}{dt}\right\}\left\{\frac{1}{2\kappa^{2}}\left(1+f^{\prime}(R)\right)-m^{\prime}(R)n_{\mathrm{DM}}\right\}\,.$ (39)
By using the $e$-foldings $N$ defined by $a=\mathrm{e}^{N}$ by assuming $a=1$ in the present Universe, we find $n_{\mathrm{DM}}=n_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}$, where $n_{\mathrm{DM\,(0)}}$ is the number density of the dark matter particles in the present Universe. In terms of $e$-foldings $N$, Eq. (IV) can be rewritten as,
$0=$ $\,\frac{1}{2\kappa^{2}}\left(R+f(R)\right)-m(R)n_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}$
+$\,+\left\{3\left(HH^{\prime}+H^{2}\right)+3H^{2}\frac{d}{dN}\right\}\left\{\frac{1}{2\kappa^{2}}\left(1+f^{\prime}(R)\right)-m^{\prime}(R)n_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}\right\}\,,$ (40)
where $H^{\prime}\equiv\frac{dH}{dN}$ and we used the relation $\frac{d}{dt}=H\frac{d}{dN}$. We assume that the time evolution of the Universe is given by a function $H=H(N)$ and therefore the scalar curvature $R$ is a function of $N$ because $R=12H(N)^{2}+6H(N)H^{\prime}(N)$. We may solve $R$ with respect to $N$, $N=N(R)$. We also assume the $N$-dependence of $\Phi=\Phi(N)\equiv\frac{1}{2\kappa^{2}}\left(1+f^{\prime}(R)\right)-m^{\prime}(R)n_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}$. Then Eq. (IV) gives
$\Psi=$ $\,\Psi(N)\equiv\frac{1}{2\kappa^{2}}\left(R+f(R)\right)-m(R)n_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}$
$=$ $\,-\left\{3\left(HH^{\prime}+H^{2}\right)\Phi(N)+3H^{2}\frac{d\Phi(N)}{dN}\right\}\,,$ (41)
Because,
$\Psi^{\prime}=\left(24HH^{\prime}+6{H^{\prime}}^{2}+6HH^{\prime\prime}\right)\Phi+3m(R)n_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}\,,$ (42)
we find,
$m\left(R\left(N\right)\right)=\frac{\mathrm{e}^{3N}}{3n_{\mathrm{DM\,(0)}}}\left(\Psi^{\prime}-\left(24HH^{\prime}+6{H^{\prime}}^{2}+6HH^{\prime\prime}\right)\Phi\right)\,,$ (43)
Using the expression of $N=N(R)$, we find $m(R)$ as a function of $R$. Then by the definition of $\Psi$, we find,
$f(R)=-R+2\kappa^{2}\Psi\left(N\left(R\right)\right)+2\kappa^{2}m\left(R\right)\mathrm{e}^{-3N(R)}\,.$ (44)
Then we find $f(R)$ as a function of $R$. First, we consider a very trivial example,
$H=H_{0}\,,\quad\Phi=\Phi_{0}\,,$ (45)
with constants $H_{0}$ and $\Phi_{0}$ to check the consistency in the above formulation. The spacetime is, of course, the de Sitter spacetime. Then Eqs. (IV), (43), and (44) tell,
$\Psi=-3{H_{0}}^{2}\Phi_{0}\,,\quad m(R)=0\,,\quad f(R)=-12{H_{0}}^{2}-6\kappa^{2}\Psi_{0}\,.$ (46)
Then we obtained the model with a cosmological constant, where $f(R)$ is a constant and there is no matter $m(R)=0$, as expected. Anyway, we find the above formulation using Eqs. (IV), (43), and (44) may work. Second, as a less non-trivial example, we consider the case that
$H^{2}={H_{0}}^{2}+{H_{1}}^{2}\mathrm{e}^{-3N}\,,$ (47)
with constants $H_{0}$ and $H_{1}$. Then we obtain $R=12H^{2}+6HH^{\prime}=12\left({H_{0}}^{2}+{H_{1}}^{2}\mathrm{e}^{-3N}\right)-9{H_{1}}^{2}\mathrm{e}^{-3N}=12{H_{0}}^{2}+3{H_{1}}^{2}\mathrm{e}^{-3N}$ and we find $N=-\frac{1}{3}\ln\left(\frac{R-12{H_{0}}^{2}}{3{H_{1}}^{2}}\right)$. Again, by using Eqs. (IV), (43), and (44) and assuming $\Phi=\Phi_{0}$ with a constant $\Phi_{0}$, we find,
$\Psi=$ $\,\left(-3{H_{0}}^{2}+\frac{3}{2}{H_{1}}^{2}\mathrm{e}^{-3N}\right)\Phi_{0}\,,\quad m\left(R\right)=\frac{3{H_{1}}^{2}\Phi_{0}}{2n_{\mathrm{DM\,(0)}}}\,,$
$f(R)=$ $\,\left(-1+\frac{2\kappa^{2}\Phi_{0}}{n_{\mathrm{DM\,(0)}}}\right)R+6\kappa^{2}{H_{0}}^{2}\Phi_{0}\left(-1+\frac{4}{n_{\mathrm{DM\,(0)}}}\right)\,.$ (48)
Therefore, we obtain Einstein’s gravity with a cosmological constant consistently, although the gravitational coupling $\kappa$ is effectively modified as $\kappa^{2}\to\frac{n_{\mathrm{DM\,(0)}}}{\Phi_{0}}$ because $F(R)=R+f(R)=\frac{2\kappa^{2}\Phi_{0}}{n_{\mathrm{DM\,(0)}}}R+6\kappa^{2}{H_{0}}^{2}\Phi_{0}\left(-1+\frac{4}{n_{\mathrm{DM\,(0)}}}\right)$. We now consider a more realistic and non-trivial model. As we found, the inverse phantom crossing occurs when the first derivative of the effective DE density $\rho^{\mathrm{eff}}_{\mathrm{DE}}$ vanishes and its second derivative is negative,
$\dot{\rho}^{\mathrm{eff}}_{\mathrm{DE}}=0\,,\quad\ddot{\rho}^{\mathrm{eff}}_{\mathrm{DE}}<0\,.$ (49)
The $e$-foldings $N$ is a monotonically increasing function of the cosmic time $t$, and the above condition (49) can be rewritten in terms of $N$, as follows,
${\rho^{\mathrm{eff}}_{\mathrm{DE}}}^{\prime}=0\,,\quad{\rho^{\mathrm{eff}}_{\mathrm{DE}}}^{\prime\prime}<0\,.$ (50)
As an example, we consider the following model,
$\rho^{\mathrm{eff}}_{\mathrm{DE}}=\Lambda_{0}+\frac{\Lambda_{1}}{\left(\mathrm{e}^{-3N}-\mathrm{e}^{-3N_{0}}\right)^{2}+C^{2}}\,.$ (51)
Here $\Lambda_{0}$ and $\Lambda_{1}$ are constants with the dimension of $[m]^{4}$, and $C$ is a dimensionless constant. The effective energy density $\rho^{\mathrm{eff}}_{\mathrm{DE}}$ has a peak at $N=N_{0}$, which corresponds to the $e$-foldings that the inverse phantom crossing occurs.
We also assume that the effective dark matter density $\rho^{\mathrm{eff}}_{\mathrm{DM}}$ behaves as $\rho^{\mathrm{eff}}_{\mathrm{DM}}\propto a^{-3}=\mathrm{e}^{-3N}$,
$\rho^{\mathrm{eff}}_{\mathrm{DM}}=\rho_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}\,.$ (52)
Then the Friedmann equation,
$\frac{3}{\kappa^{2}}H^{2}=\rho^{\mathrm{eff}}_{\mathrm{DE}}+\rho^{\mathrm{eff}}_{\mathrm{DM}}\,,$ (53)
indicates that,
$H^{2}=\frac{3}{\kappa^{2}}\left\{\Lambda_{0}+\frac{\Lambda_{1}}{\left(\mathrm{e}^{-3N}-\mathrm{e}^{-3N_{0}}\right)^{2}+C^{2}}+\rho_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}\right\}\,.$ (54)
Therefore, the scalar curvature $R$ is given by,
$R=$  $\,12H^{2}+6HH^{\prime}$
$=$ $\,\frac{36}{\kappa^{2}}\left\{\Lambda_{0}+\frac{\Lambda_{1}}{\left(\mathrm{e}^{-3N}-\mathrm{e}^{-3N_{0}}\right)^{2}+C^{2}}+\rho_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}\right\}$
$\,+\frac{9}{\kappa^{2}}\left\{\frac{6\Lambda_{1}\left(\mathrm{e}^{-3N}-\mathrm{e}^{-3N_{0}}\right)\mathrm{e}^{-3N}}{\left\{\left(\mathrm{e}^{-3N}-\mathrm{e}^{-3N_{0}}\right)^{2}+C^{2}\right\}^{2}}-3\rho_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}\right\}\,.$ (55)
In principle, Eq. (IV) can be solved with respect to $N$, $N=N(R)$, although it is difficult to do it explicitly. The phantom crossing occurs when $N\sim N_{0}$, given by,
$R=$ $\,R_{0}+R_{1}\left(N-N_{0}\right)+\mathcal{O}\left(\left(N-N_{0}\right)^{2}\right)\,,$
$R_{0}\equiv\frac{9}{\kappa^{2}}\left(4\Lambda_{0}+\frac{4\Lambda_{1}}{C^{2}}+\rho_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N_{0}}\right)\,,\quad R_{1}\equiv-\frac{27}{\kappa^{2}}\left(\frac{30\Lambda_{1}\mathrm{e}^{-3N_{0}}}{C^{4}}+\rho_{\mathrm{DM\,(0)}}\right)\,,$ (56)
which can be solved with respect to $N$,
$N=N_{0}+\frac{R-R_{0}}{R_{1}}+\mathcal{O}\left(\left(R-R_{0}\right)^{2}\right)\,.$ (57)
In principle, by using Eqs. (IV), (43), and (44) and properly assuming $\Phi=\Phi_{0}$, we may find the form of $f(R)$ and $m(R)$ although it is very difficult to find their explicit forms.
V Oscillating Hubble Rate
As in Elizalde:2011ds ; Elizalde:2010ts , we consider the model where the Hubble rate $H$ oscillates and therefore the Universe iterates the phantom crossing and the inverse phantom crossing. In this section, we consider such an oscillation for the total energy density or the behaviour of $H$. A simple model of this sort is given by,
$H^{2}={H_{0}}^{2}\left(1+C\sin\left(\omega N\right)\right)\,,$ (58)
where $H_{0}$, $C$, and $\omega$ are constants, and we assume $10<C<$ 1 and $\omega>0$. Then we find,
$HH^{\prime}=\frac{{H_{0}}^{2}C\omega}{2}\cos\left(\omega N\right)\,.$ (59)
Because we are considering the expanding Universe, where $H>0$, when $-\frac{\pi}{2}+2n\pi<\omega N<\frac{\pi}{2}+2n\pi$ with an integer $n$, $H^{\prime}$ is positive and therefore, the Universe is in the phantom phase. On the other hand, when $\frac{\pi}{2}+2n\pi<\omega N<\frac{3\pi}{2}+2n\pi$ with an integer $n$, $H^{\prime}$ is negative and therefore, the Universe is in the non-phantom phase. The phantom crossing occurs when $\omega N=-\frac{\pi}{2}+2n\pi$ with an integer $n$ and the inverse phantom crossing occurs when $\omega N=\frac{\pi}{2}+2n\pi$ with an integer $n$. Under the assumption (58), the scalar curvature $R$ is given by
$R=12H^{2}+6HH^{\prime}=3{H_{0}}^{2}\left(4+\sqrt{16+\omega^{2}}C\sin\left(\omega N+\theta_{0}\right)\right)\,.$ (60)
Here $\theta_{0}$ is defined by $\tan\theta_{0}=\frac{\omega}{4}$. Eq. (60) can be solved with respect to $N$,
$N=-\frac{\theta_{0}}{\omega}\sin^{-1}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\,.$ (61)
Here $sin^{-1}$ is the inverse function of $\sin$ function. The $e$-foldings $N$ is explicitly given as a function of $R$. Furthermore, we find,
$\sin\left(\omega N\right)=$ $\,\frac{4}{\sqrt{16+\omega^{2}}}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)-\frac{\omega}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}\,,$
$\cos\left(\omega N\right)=$ $\,\frac{4}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}+\frac{\omega}{\sqrt{16+\omega^{2}}}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\,,$ (62)
which we will use shortly. Again, we use Eqs. (IV), (43), and (44). Then by assuming $\Phi=\Phi_{0}$, we find,
$\Psi(N)=$ $\,-3\Phi_{0}{H_{0}}^{2}\left(1+C\sin\left(\omega N\right)+\frac{C\omega}{2}\cos\left(\omega N\right)\right)\,,$
$m(N)=$ $\,-\frac{\Phi_{0}{H_{0}}^{2}C\mathrm{e}^{3N}}{n_{\mathrm{DM\,(0)}}}\left\{\frac{9\omega^{2}}{2}\sin\left(\omega N\right)-15\omega\cos\left(\omega N\right)\right\}\,,$
$f=$ $\,-R-2\kappa^{2}\Phi_{0}{H_{0}}^{2}\left[3+C\left\{\left(3+\frac{9\omega^{2}}{2n_{\mathrm{DM\,(0)}}}\right)\sin\left(\omega N\right)\right.\right.$
$\,\left.\left.+\left(\frac{3}{2}-\frac{15}{n_{\mathrm{DM\,(0)}}}\right)\omega\cos\left(\omega N\right)\right\}\right]\,.$ (63)
Then by using (61), we find the $R$ dependence of $m=m(R)$ and $f=f(R)$,
$m(R)=$ $\,-\frac{\Phi_{0}{H_{0}}^{2}C\mathrm{e}^{-\frac{3\theta_{0}}{\omega}\sin^{-1}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)}}{n_{\mathrm{DM\,(0)}}\sqrt{16+\omega^{2}}}\left\{3\omega^{2}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\right.$
$\,\left.-\left(\frac{9\omega^{3}}{2}+60\omega\right)\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}\right\}\,,$
$f(R)=$ $\,-R-2\kappa^{2}\Phi_{0}{H_{0}}^{2}\left[3+\frac{C}{\sqrt{16+\omega^{2}}}\left\{\left(12+\frac{3}{2}\omega^{2}+\frac{3\omega^{2}}{n_{\mathrm{DM\,(0)}}}\right)\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\right.\right.$
$\,\left.\left.+\left(3\omega-\frac{120\omega+9\omega^{3}}{2n_{\mathrm{DM\,(0)}}}\right)\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}\right\}\right]$
In this model, the mass $m=m(R)$ of the dark matter depends on the scalar curvature $R$ non-trivially.
VI Scalaron as Dark Matter with Curvature-dependent Mass
There is a scenario in which the dark matter could be a scalar mode coming from the $F(R)$ gravity Nojiri:2008nt ; Katsuragawa:2016yir ; Katsuragawa:2017wge . In this scenario, the mass of the scalar mode naturally depends on the curvature. Then the particles corresponding to the scalar mode might become dark matter, which may explain the DESI observation.
The action of the $F(R)$ gravity is given by replacing the scalar curvature $R$ in the Einstein-Hilbert action, which is,
$S_{\mathrm{EH}}=\int d^{4}x\sqrt{-g}\left(\frac{R}{2\kappa^{2}}+\mathcal{L}_{\mathrm{matter}}\right)\,,$ (65)
by using some appropriate function of the scalar curvature, as follows,
$S_{F(R)}=\int d^{4}x\sqrt{-g}\left(\frac{F(R)}{2\kappa^{2}}+\mathcal{L}_{\mathrm{matter}}\right)\,.$ (66)
In Eqs. (65) and (66), $\mathcal{L}_{\mathrm{matter}}$ is the matter Lagrangian density.
By varying the action (66) with respect to the metric, we obtain the equation of motion for the $F(R)$ gravity theory as follows,
$\frac{1}{2}g_{\mu\nu}F(R)-R_{\mu\nu}F^{\prime}(R)-g_{\mu\nu}\Box F^{\prime}(R)+\nabla_{\mu}\nabla_{\nu}F^{\prime}(R)=-\frac{\kappa^{2}}{2}T_{\mathrm{matter}\,\mu\nu}\,.$ (67)
For a spatially flat FLRW Universe, Eq. (67) gives the field equations,
$0=$ $\,-\frac{F(R)}{2}+3\left(H^{2}+\dot{H}\right)F^{\prime}(R)-18\left(4H^{2}\dot{H}+H\ddot{H}\right)F^{\prime\prime}(R)+\kappa^{2}\rho_{\mathrm{matter}}\,,$ (68)
$0=$ $\,\frac{F(R)}{2}-\left(\dot{H}+3H^{2}\right)F^{\prime}(R)+6\left(8H^{2}\dot{H}+4{\dot{H}}^{2}+6H\ddot{H}+\dddot{H}\right)F^{\prime\prime}(R)$
$\,+36\left(4H\dot{H}+\ddot{H}\right)^{2}F^{\prime\prime\prime}(R)+\kappa^{2}p_{\mathrm{matter}}\,,$ (69)
where the Hubble rate $H$ is equal to $H=\dot{a}/a$. In terms of the Hubble rate $H$, the scalar curvature $R$ is equal to $R=12H^{2}+6\dot{H}$. We now consider the scalar-tensor representation of the $F(R)$ gravity. We introduce an auxiliary field $A$ and rewrite the action (66) of the $F(R)$ gravity in the following form:
$S=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}\left\{F^{\prime}(A)\left(R-A\right)+F(A)\right\}\,.$ (70)
We obtain $A=R$ by the variation of the action with respect to $A$ and by substituting the obtained equation $A=R$ into the action (70), we find that the action in (66) is reproduced. If we rescale the metric by a kind of scale transformation,
$g_{\mu\nu}\to\mathrm{e}^{\sigma}g_{\mu\nu}\,,\quad\sigma=-\ln F^{\prime}(A)\,,$ (71)
we obtain the action in the Einstein frame,
$S_{E}=$ $\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}\left(R-\frac{3}{2}g^{\rho\sigma}\partial_{\rho}\sigma\partial_{\sigma}\sigma-V(\sigma)\right)\,,$
$V(\sigma)=$ $\mathrm{e}^{\sigma}g\left(\mathrm{e}^{-\sigma}\right)-\mathrm{e}^{2\sigma}f\left(g\left(\mathrm{e}^{-\sigma}\right)\right)=\frac{A}{F^{\prime}(A)}-\frac{F(A)}{F^{\prime}(A)^{2}}\,.$ (72)
Here $g\left(\mathrm{e}^{-\sigma}\right)$ is given by solving the equation $\sigma=-\ln\left(1+f^{\prime}(A)\right)=-\ln F^{\prime}(A)$ as $A=g\left(\mathrm{e}^{-\sigma}\right)$. Due to the scale transformation (71), a coupling of the scalar field $\sigma$ with usual matter is introduced. The mass of the scalar field $\sigma$ is given by,
${m_{\sigma}}^{2}\equiv\frac{3}{2}\frac{d^{2}V(\sigma)}{d\sigma^{2}}=\frac{3}{2}\left\{\frac{A}{F^{\prime}(A)}-\frac{4F(A)}{\left(F^{\prime}(A)\right)^{2}}+\frac{1}{F^{\prime\prime}(A)}\right\}\,.$ (73)
Because $A=R$, the mass depends on the curvature. In the case of Einstein’s gravity, we find $F^{\prime\prime}(A)=0$ and the mass diverges, as expected. The scalar mode decouples when $F^{\prime\prime}(A)=0$.
The scenario that the $F(R)$ scalaron-chameleon can be dark matter has been investigated Nojiri:2008nt ; Katsuragawa:2016yir ; Katsuragawa:2017wge . If the mass increases when the curvature $R$ decreases, the problem of the DESI observation might be solved. As an example, we may consider the following simple model
$F(R)=R-\Lambda+\alpha R^{2}\,.$ (74)
Then (73) tells,
${m_{\sigma}}^{2}=\frac{3}{2}\left\{\frac{R}{1+2\alpha R}-\frac{4\left(R-\Lambda+\alpha R^{2}\right)}{\left(1+2\alpha R\right)^{2}}+\frac{1}{2\alpha}\right\}\,.$ (75)
When $R$ is large ${m_{\sigma}}^{2}$ goes to vanish and when $R\to 0$, we find ${m_{\sigma}}^{2}\to 4\Lambda+\frac{1}{2\alpha}$. Therefore, the mass increases when $R$ decreases.
We now rewrite Eq. (75) as follows,
${m_{\sigma}}^{2}=$ $\,\frac{3}{2}\frac{\left(4-8+4\right)\alpha^{2}R^{2}+\left(2-8+4\right)\alpha R+8\alpha\Lambda+1}{2\alpha\left(1+2\alpha R\right)^{2}}$
$=$ $\,\frac{3\left(-2\alpha R+8\alpha\Lambda+1\right)}{4\alpha\left(1+2\alpha R\right)^{2}}\,.$ (76)
Therefore, when $R>4\Lambda+\frac{1}{2\alpha}$, the scalar mode becomes tachyonic. This is not surprising because it corresponds to the slow-roll during inflation. The model (74) could describe inflation due to the last term, $\alpha R^{2}$, when $R$ is large. The second term could play the role of a small cosmological term for the late-time accelerated expansion, when $R$ is small. Then we expect $\alpha\sim\left(10^{13}\,\mathrm{GeV}\right)^{-2}$ and $\Lambda\sim\left(10^{-33}\,\mathrm{eV}\right)^{2}$. Therefore, we may estimate $m_{\sigma}$ as $m_{\sigma}\sim 10^{13}\,\mathrm{GeV}$, which is large enough. Anyway, the model (74) simply describes both the inflation and the DE.
VII Towards a Realistic Model: A Theoretical Approach
In order to obtain a realistic model, we need to clarify the creation of the scalaron particles and to check the consistency of the parameters. First, we should note that, as is clear from the scale transformation (71), there is an interaction between the scalaron and the matter, including the standard model particles. Then, in the thermal equilibrium, the scalaron could be created by the interaction. Because the coupling is given by the gravitational coupling, which is the inverse of the Planck mass, the interaction is very weak, and therefore, a sufficient amount of the scalaron particles could be created, and the mass of the scalaron must be small enough. Eq. (75) could tell that the mass might be small just after the inflationary era. There might also be gravitational particle production.
We should note that the DESI observation corresponds to the redshift region $3.50<z<$ 3.5. If we choose the parameter $\alpha$ in the model in (74) to be $\alpha\sim\left(10^{13}\,\mathrm{GeV}\right)^{-2}$ so that the model could also describe the inflation, Eq. (VI) tells that in the DESI redshift, the mass is already heavy enough, ${m_{\sigma}}^{2}\sim\frac{3}{4\alpha}$, and therefore the inverse phantom crossing could not occur. In order to avoid this problem, we may choose $1/\alpha$ to be small enough. In the laboratory, the energy density of the redshift region $3.50<z<$ 3.5 can be easily realised, and the fifth force experiment excludes such a light mode. The coupling between the scalaron and the matter shifts the mass due to the Chameleon mechanism, but the scalaron could be light enough to be excluded from the experiments. In order to avoid the problem, we may modify the model so that $\alpha$ could depend on the scalar curvature. The problem might be solved if $1/\alpha$ is large enough, even in the redshift region $3.50<z<$ 3.5 to be consistent with the fifth force experiment, and after that $1/\alpha$ further increases to generate the apparent phantom crossing.
VIII Some Explicit Examples of Oscillating DE with Phantom Crossing
Now let us consider some realistic models of $F(R)$ gravity DE which we will extensively study using numerical methods, and we will obtain the cosmological parameters of these models. We will be mainly interested in the DE density and DE EoS parameter, which are constrained by the data, and we will numerically show that the models exhibit late-time EoS oscillations and a phantom crossing. Let us consider the $F(R)$ gravity action in the presence of perfect matter fluids in (66). Again we assume that the $F(R)$ gravity has the form of (4) and upon varying the action (66) with respect to the FLRW metric, we obtain,
$3F_{R}H^{2}=$ $\,\kappa^{2}\rho_{\mathrm{matter}}+\frac{F_{R}R-F}{2}-3H\dot{F}_{R}\,,$ (77)
$-2F_{R}\dot{H}=$ $\,\kappa^{2}(\rho_{\mathrm{matter}}+p_{\mathrm{matter}})+\ddot{F}-H\dot{F}\,,$ (78)
with $F_{R}=\frac{\partial F}{\partial R}$ and the “dot” indicates differentiation with respect to cosmic time, and $\rho_{\mathrm{matter}}$ and $p_{\mathrm{matter}}$ stand for the matter fluids energy density and pressure. We can rewrite the field equations (77), (78) in the Einstein-Hilbert gravity form,
$3H^{2}=$ $\,\kappa^{2}\rho_{\mathrm{total}}\,,$ (79)
$-2\dot{H}=$ $\,\kappa^{2}(\rho_{\mathrm{total}}+p_{\mathrm{total}})\,,$ (80)
where $\rho_{\mathrm{total}}$ stands for the total energy density of the total effective cosmological fluid, which includes the $F(R)$ gravity contribution, and $p_{\mathrm{total}}$ denotes the total pressure. The total cosmological fluid is composed of three parts: the cold dark matter one ($\rho_{\mathrm{matter}}$), the radiation ($\rho_{\mathrm{r}}$) and the geometric part of $F(R)$ gravity ($\rho_{\mathrm{DE}}$). Therefore, we have, $\rho_{\mathrm{total}}=\rho_{\mathrm{matter}}+\rho_{\mathrm{r}}+\rho_{\mathrm{DE}}$ and in addition $p_{\mathrm{total}}=p_{\mathrm{matter}}+p_{\mathrm{r}}+p_{\mathrm{DE}}$. It is the geometric fluid which drives the dynamical evolution in the late-time era, and with its energy density and pressure being,
$\rho_{\mathrm{DE}}=$ $\,\frac{F_{R}R-F}{2}+3H^{2}(1-F_{R})-3H\dot{F}_{R}\,,$ (81)
$p_{\mathrm{DE}}=$ $\,\ddot{F}-H\dot{F}+2\dot{H}(F_{R}-1)-\rho_{\mathrm{DE}}\,.$ (82)
In the following, instead of the cosmic time, we shall use the redshift as a dynamical variable, defined as,
$1+z=\frac{1}{a}\,,$ (83)
and in addition, we introduce the statefinder parameter $y_{H}(z)$ Hu:2007nk ; Bamba:2012qi ; Nojiri:2017ncd ,
$y_{H}(z)=\frac{\rho_{\mathrm{DE}}}{\rho_{\mathrm{matter}}^{(0)}}=\frac{H^{2}}{{m_{s}}^{2}}-(1+z)^{3}-\chi(1+z)^{4}\,,$ (84)
with $\rho_{\mathrm{matter}}^{(0)}$ denoting the energy density of the cold dark matter at the present epoch, and in addition ${m_{s}}^{2}=\frac{\kappa^{2}\rho_{\mathrm{matter}}^{(0)}}{3}={H_{0}}^{2}\Omega_{\mathrm{matter}}=1.37\times 10^{-67}\,\mathrm{eV}^{2}$ while $\chi$ is $\chi=\frac{\rho_{\mathrm{r}}^{(0)}}{\rho_{\mathrm{matter}}^{(0)}}\simeq 3.1\times 10^{-4}$, with $\rho_{\mathrm{r}}^{(0)}$ denoting the present epoch radiation energy density. Combining Eqs. (4), (79), and (84), we can rewrite the Friedmann equation in terms of the statefinder $y_{H}$ in the following way,
$\frac{d^{2}y_{H}}{dz^{2}}+J_{1}\frac{dy_{H}}{dz}+J_{2}y_{H}+J_{3}=0\,,$ (85)
where the dimensionless functions $J_{1}$ , $J_{2}$ , $J_{3}$ are defined as,
$J_{1}=$ $\,\frac{1}{(z+1)}\left(-3-\frac{1}{y_{H}+(z+1)^{3}+\chi(z+1)^{4}}\frac{1-F_{R}}{6{m_{s}}^{2}F_{RR}}\right)\,,$ (86)
$J_{2}=$ $\,\frac{1}{(z+1)^{2}}\left(\frac{1}{y_{H}+(z+1)^{3}+\chi(z+1)^{4}}\frac{2-F_{R}}{3{m_{s}}^{2}F_{RR}}\right)\,,$ (87)
$J_{3}=$ $\,-3(z+1)-\frac{(1-F_{R})\left((z+1)^{3}+2\chi(z+1)^{4}\right)+(R-F)/\left(3{m_{s}}^{2}\right)}{(z+1)^{2}\left(y_{H}+(z+1)^{3}+\chi(z+1)^{4}\right)}\frac{1}{6{m_{s}}^{2}F_{RR}}\,,$ (88)
and furthermore $F_{RR}=\frac{\partial^{2}F}{\partial R^{2}}$. Moreover, the Ricci scalar is,
$R=12H^{2}-6HH_{z}(1+z)\,,$ (89)
and in terms of $y_{H}$, we obtain,
$R(z)=3\,{m_{s}}^{2}\left(-(z+1)\,\frac{dy_{H}(z)}{dz}+4y_{H}(z)+(1+z)^{3}\right)\,.$ (90)
We will solve Eq. (85) numerically for the late-time redshift interval $z=[0,10]$, using the following initial conditions at the final redshift $z_{f}=10$ Bamba:2012qi ,
$y_{H}(z_{f})=\frac{\Lambda}{3{m_{s}}^{2}}\left(1+\frac{1+z_{f}}{1000}\right)\,,\quad\left.\frac{dy_{H}(z)}{dz}\right|_{z=z_{f}}=\frac{1}{1000}\frac{\Lambda}{3{m_{s}}^{2}}\,,$ (91)
where $\Lambda\simeq 11.895\times 10^{-67}\,\mathrm{eV}^{2}$. We can express the physical cosmological quantities using the statefinder $y_{H}(z)$, and we have,
$H(z)=m_{s}\sqrt{y_{H}(z)+(1+z)^{3}+\chi(1+z)^{4}}\,.$ (92)
and the Ricci scalar is,
$R(z)=3{m_{s}}^{2}\left(4y_{H}(z)-(z+1)\frac{dy_{H}(z)}{dz}+(z+1)^{3}\right)\,,$ (93)
while the DE density parameter $\Omega_{\mathrm{DE}}(z)$ is,
$\Omega_{\mathrm{DE}}(z)=\frac{y_{H}(z)}{y_{H}(z)+(z+1)^{3}+\chi(z+1)^{4}}\,,$ (94)
and the DE EoS parameter is,
$w_{\mathrm{DE}}(z)=-1+\frac{1}{3}(z+1)\frac{1}{y_{H}(z)}\frac{dy_{H}(z)}{dz}\,.$ (95)
Finally, the total EoS parameter is,
$w_{\mathrm{total}}(z)=\frac{2(z+1)H^{\prime}(z)}{3H(z)}-1\,.$ (96)
while the deceleration parameter is,
$q(z)=-1-\frac{\dot{H}}{H^{2}}=-1-(z+1)\frac{H^{\prime}(z)}{H(z)}\,,$ (97)
where the “prime” denotes differentiation with respect to the redshift. In addition, and for comparison reasons, let us quote here the Hubble rate for the $\Lambda$CDM model,
$H_{\Lambda}(z)=H_{0}\sqrt{\Omega_{\Lambda}+\Omega_{\mathrm{matter}}(z+1)^{3}+\Omega_{\mathrm{r}}(z+1)^{4}}\,,$ (98)
where $\Omega_{\Lambda}\simeq 0.68136$ and $\Omega_{\mathrm{matter}}\simeq 0.3153$ and $H_{0}\simeq 1.37187\times 10^{-33}\,\mathrm{eV}$ according to the latest 2018 Planck data Planck:2018vyg . Now let us focus on the phenomenology of two interesting models which exhibit remarkable behaviour at late times. We start off with the following model,
$F(R)=R+\frac{R^{2}}{M^{2}}-\frac{b}{c+\exp(-R/R_{0})}\,,$ (99)
with $b$ and $R_{0}$ are free parameters with mass dimensions $[m]^{2}$ and also $c$ is a dimensionless parameter. We obtained a viable phenomenology for the choice $b=20\Lambda,$ $c=2$, and $R_{0}={m_{s}}^{2}/0.00091$. In Fig. 1 we plot the deceleration parameter for this model (red curve) compared with the $\Lambda$CDM (black curve), and the DE EoS for various redshift ranges. As can be seen, the DE EoS is strongly oscillating at late times, and the model ends up in a quintessential regime. The DE energy density parameter and the DE EoS parameter at $z=0$ are for this model $\Omega_{\mathrm{DE}}=0.6918$ and $w_{\mathrm{DE}}=-0.9974$ which comply with the latest Planck constraints.
Figure 1: Plots of the deceleration parameter $q(z)$ (upper left plot) and the DE EoS parameter $w_{\mathrm{DE}}(z)$ (right and bottom plot) as functions of the redshift for the model Eq. (99) for $b=20\Lambda$, $c=2$, and $R_{0}={m_{s}}^{2}/0.00091$.
Another phenomenologically interesting $F(R)$ model is,
$F(R)=R+\frac{R^{2}}{M^{2}}-\alpha\frac{b(R/R_{0})^{n}}{c(R/R_{0})^{n}+d}\,,$ (100)
with $b$, $c$, $d$, $n$ being dimensionless parameters and $\alpha$, $R_{0}$ are also free parameters with mass dimensions $[m]^{2}$. With the choice $\alpha=1.4\Lambda$, $b=1$, $c=0.2$, $d=0.04$, $R_{0}={m_{s}}^{2}$, and $n=0.3$, we obtain an interesting DE oscillating behaviour which is also phenomenologically viable. Specifically, in Fig. 2, we plot the deceleration parameter (red curve) compared with the $\Lambda$CDM (black curve), and also the DE EoS for various redshift ranges, focusing on small redshifts. As can be seen in this case, too, the DE EoS is strongly oscillating at late times and at small redshifts before the present day, and the model ends up in a quintessential regime in this case, too. The DE energy density parameter and the DE EoS parameter at $z=0$ for this model are $\Omega_{\mathrm{DE}}=0.6851$ and $w_{\mathrm{DE}}=-0.9887$ which are again compatible with the latest Planck constraints.
Figure 2: Plots of the deceleration parameter $q(z)$ (upper left plot) and the DE EoS parameter $w_{\mathrm{DE}}(z)$ (right and bottom plot) as functions of the redshift for the model of Eq. (100) with $\alpha=1.4\Lambda$, $b=1$, $c=0.2$, $d=0.04$, $R_{0}={m_{s}}^{2}$, and $n=0.3$.
Thus in this section, using a numerical approach, we demonstrated that it is possible to obtain oscillating DE eras using rather simple $F(R)$ gravity models. Note that the models presented in this section were also extensively examined regarding their phenomenology in Ref. Odintsov:2025jfq and the model (99) is found to be fully compatible with the observational data, including the latest DESI data.
IX Conclusions
In this paper, we studied the possibility of having dynamical DE which crosses the phantom divide line and oscillates at late times. We examined this problem both theoretically and numerically and proved that $F(R)$ gravity provides a consistent phenomenological framework for naturally describing the DE era without resorting to phantom fields. Specifically, we discussed the inverse-phantom crossing in $F(R)$ gravity from a theoretical viewpoint and discussed the conditions that the $F(R)$ gravity fluid must satisfy in order to achieve the inverse phantom crossing. We also discussed the apparent phantom crossing issue, further extending the considerations made. Furthermore, we developed the theoretical study of several models that may exhibit phantom crossing and discussed from a theoretical viewpoint these models and their DE behaviour, in a qualitative way, though. Moreover, several oscillating forms of the Hubble rate were considered, in terms of the $e$-foldings number and their realisation in $F(R)$ gravity was also considered. In addition, we considered the effects of a scalaron as dark matter with a curvature-dependent mass on the DE era. We also studied in a quantitative way the predictions on the DE era of two explicit $F(R)$ gravity models, one of which is also compatible with the DESI data, apart from the Planck data. We studied the behaviour of the deceleration parameter and the DE EoS, and showed that it is possible to obtain a phantom crossing EoS and, in addition, an oscillating DE EoS at late times. Thus, we proved that indeed $F(R)$ gravity offers a consistent framework which can generate a phantom to quintessence transition in a natural way without phantom fields, and moreover, some of the models realise DE oscillations, again built in the model, not artificially added. Now the question is why should $F(R)$ gravity should be considered a viable extension of Einstein’s gravity. The answer is simple: Einstein’s gravitational Lagrangian density is described simply by $R$, so an Occam’s razor approach is that an $F(R)$ gravity extends it. Of course, someone may also seek if such oscillating and phantom divide crossing phenomena occur in other gravities, like Einstein-Gauss-Bonnet theories, but the natural extension of Einstein’s gravity is $F(R)$ gravity and it enjoys an elevated role among all modified gravity models.
Acknowledgements
We are indebted to Taishi Katsuragawa for discussions. This work was partially supported by the program Unidad de Excelencia Maria de Maeztu CEX2020-001058-M, Spain (S.D.O).

### v2

@@
 Abstract
-In this work, we shall consider how a dynamical oscillating and phantom crossing dark energy era can be realised in the context of $F(R)$ gravity.
-We approach the topic from a theoretical standpoint, considering all the conditions that may lead to a consistent phantom crossing behaviour and separately how the $F(R)$ gravity context may realise an oscillating dark energy era.
-Apart from our qualitative considerations, we study quantitatively two $F(R)$ gravity dark energy models, which are viable cosmologically and also exhibit simultaneously phantom crossing behaviour and oscillating dark energy.
-We consider these models by numerically solving the field equations using appropriate statefinder parameters engineered for dark energy studies.
-As we show, $F(R)$ provides a natural extension of Einstein’s general relativity, which can naturally realise a transition from a phantom era to a quintessential era, a feature supported by recent observational data, without resorting to phantom scalar fields to realise the phantom evolution.
-preprint: KEK-TH-2734, KEK-Cosmo-0383
+In this work, we shall consider how a dynamical oscillating and phantom crossing dark energy era can be realized in the context of $F(R)$ gravity.
+We approach the topic from a theoretical standpoint considering all the conditions that may lead to a consistent phantom crossing behavior and separately how the $F(R)$ gravity context may realize oscillating dark energy era.
+Apart from our qualitative considerations, we study in a quantitative way two $F(R)$ gravity dark energy models which are viable cosmologically and also exhibit simultaneously phantom crossing behavior and also oscillating dark energy.
+We consider these models by solving numerically the field equations using appropriate statefinder parameters engineered for dark energy studies.
+As we show, $F(R)$ provides a natural extension of Einstein’s general relativity which can naturally realize a transition from a phantom era to a quintessential era, a feature supported by recent observational data, without resorting to phantom scalar fields to realize the phantom evolution.
 I Introduction
-The observation of the late-time acceleration of the Universe, known as dark energy (DE) era is one of the most unexpected features of our world.
-In the context of simple general relativity, the DE era can be generated by quintessence scalar fields, however, the latest observations of the Dark Energy Spectroscopic Instrument (DESI) DESI:2024mwx ; DESI:2025zgx indicate that the DE was in the phantom regime in the near past.
+The observation of the late-time acceleration of the Universe, known as dark energy (DE) era is one of the most unexpected features of our Universe.
+In the context of simple general relativity, the DE era can be generated by quintessence scalar fields, however the latest observations of the Dark Energy Spectroscopic Instrument (DESI) DESI:2024mwx ; DESI:2025zgx indicate that the DE was in the phantom regime in the near past.
 This feature can be described in the context of general relativity only by using phantom scalar fields, which is a rather unattractive feature.
 Also and more importantly, the DESI data indicate a transition from a phantom to a quintessential regime for DE.
-This is questionable if it can be realised in the context of general relativity.
-There are many ways to produce results compatible with the DESI data, for example Odintsov:2024woi ; Dai:2020rfo ; He:2020zns ; Nakai:2020oit ; DiValentino:2020naf ; Agrawal:2019dlm ; Ye:2020btb ; Vagnozzi:2021tjv ; Desmond:2019ygn ; Hogas:2023pjz ; OColgain:2018czj ; Vagnozzi:2019ezj ; Krishnan:2020obg ; Colgain:2019joh ; Vagnozzi:2021gjh ; Lee:2022cyh ; Krishnan:2021dyb ; Ye:2021iwa ; Ye:2022afu ; Verde:2019ivm ; Menci:2024rbq ; Adil:2023ara ; Reeves:2022aoi ; Ferlito:2022mok ; Vagnozzi:2021quy ; DiValentino:2020evt ; Sabogal:2024yha ; Giare:2024smz ; DiValentino:2025sru ; Odintsov:2025kyw ; vanderWesthuizen:2025iam ; Paliathanasis:2025dcr ; Odintsov:2025jfq ; Kessler:2025kju ; Nojiri:2025jhh ; Ozulker:2025ehg and among these frameworks, $F(R)$ gravity Nojiri:2017ncd ; Capozziello:2011et ; Nojiri:2006ri ; Nojiri:2010wj ; Nojiri:2003ft ; Capozziello:2005ku ; Hwang:2001pu ; Song:2006ej ; Faulkner:2006ub ; Olmo:2006eh ; Sawicki:2007tf ; Faraoni:2007yn ; Carloni:2007yv ; Nojiri:2007as ; Deruelle:2007pt ; Appleby:2008tv ; Dunsby:2010wg ; Oikonomou:2025qub enjoys an elevated role.
+This is questionable if it can be realized in the context of general relativity.
+There are many ways to produce results compatible with the DESI data, for example OdintsovSGS:2024 ; Dai:2020rfo ; He:2020zns ; Nakai:2020oit ; DiValentino:2020naf ; Agrawal:2019dlm ; Ye:2020btb ; Vagnozzi:2021tjv ; Desmond:2019ygn ; Hogas:2023pjz ; OColgain:2018czj ; Vagnozzi:2019ezj ; Krishnan:2020obg ; Colgain:2019joh ; Vagnozzi:2021gjh ; Lee:2022cyh ; Krishnan:2021dyb ; Ye:2021iwa ; Ye:2022afu ; Verde:2019ivm ; Menci:2024rbq ; Adil:2023ara ; Reeves:2022aoi ; Ferlito:2022mok ; Vagnozzi:2021quy ; DiValentino:2020evt ; Sabogal:2024yha ; Giare:2024smz ; DiValentino:2025sru ; Odintsov:2025kyw ; vanderWesthuizen:2025iam ; Paliathanasis:2025dcr ; Odintsov:2025jfq ; Kessler:2025kju and among these frameworks, $F(R)$ gravity reviews1 ; reviews2 ; reviews3 ; reviews4 ; Nojiri:2003ft ; Capozziello:2005ku ; Hwang:2001pu ; Song:2006ej ; Faulkner:2006ub ; Olmo:2006eh ; Sawicki:2007tf ; Faraoni:2007yn ; Carloni:2007yv ; Nojiri:2007as ; Deruelle:2007pt ; Appleby:2008tv ; Dunsby:2010wg ; Oikonomou:2025qub enjoys an elevated role.
 There are recent studies which discuss the ability of $F(R)$ gravity to be compatible with the DESI data Odintsov:2025jfq .
-Another more fascinating behaviour of DE based on the DESI data was demonstrated in Ref.
+Another more fascinating behavior of DE based on the DESI data was demonstrated in Ref.
 Kessler:2025kju , which indicates that the DE might not be only dynamical and evolving from phantom to quintessence, but it also shows strong indications of an oscillating equation of state (EoS).
-In this work, we aim to investigate whether a phantom involving and oscillating DE may be realised by $F(R)$ gravity in general.
-We approach the problem theoretically, finding the conditions which may allow such a phenomenological behaviour to be produced, and also we present several phenomenological models from a theoretical point of view.
-In addition, we study numerically the DE era of some explicit $F(R)$ gravity models which can realise a viable Planck compatible Planck:2018vyg and in some cases DESI compatible late-time behaviour.
+In this work we aim to investigate whether a phantom involving and oscillating DE may be realized by $F(R)$ gravity in general.
+We approach the problem theoretically, finding the conditions which may allow such a phenomenological behavior to be produced and also we present several phenomenological models from a theoretical point of view.
+In addition, we study numerically the DE era of some explicit $F(R)$ gravity models which can realize a viable Planck compatible Planck:2018vyg and in some cases DESI compatible late-time behavior.
 We also demonstrate that these models produce an oscillating DE era and also allow for phantom to quintessence transitions.
-This paper is organised as follows: In Section II, we discuss the inverse-phantom processing in $F(R)$ gravity from a theoretical viewpoint.
-The analysis is continued in Section III, where we discuss the apparent phantom crossing.
+This paper is organized as follows: In section II we discuss the inverse-phantom prossing in $F(R)$ gravity from a theoretical viewpoint.
+The analysis is continued in section III where we discuss the apparent phantom crossing.
 In section IV, the theoretical study of several models that exhibit phantom crossing is considered.
-The oscillating Hubble rate is considered in Section V, while in Section VI, the effects of a scalaron as dark matter with curvature-dependent mass on the DE era are considered.
-In Section VII, we discuss the features of a realistic model, while in Section VIII, we numerically study two explicit examples of oscillating $F(R)$ gravity DE with phantom crossing.
+The oscillating Hubble rate is considered in section V, while in section VI the effects of a scalaron as dark matter with curvature-dependent mass on the DE era are considered.
+In section VII, we discuss the features of a realistic model while in section VIII we numerically study two explicit examples of oscillating $F(R)$ gravity DE with phantom crossing.
 Finally, the conclusions follow at the end of the article.
@@
 Consider the vacuum $F(R)$ gravity action,
-$\mathcal{S}=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}F(R)\,,$ (1)
+$\centering\mathcal{S}=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}F(R)\,$,\@add@centering (1)
 and a flat Friedmann-Lemaître-Robertson-Walker (FLRW) metric with the following line element,
 $\centering ds^{2}=-dt^{2}+a(t)\sum_{i=1}^{3}dx_{i}^{2}\,.\@add@centering$ (2)
-Let $w$ be the DE equation of state (EoS) parameter, that is, the ratio of the pressure of the DE fluid with respect to the energy density of the DE fluid.
-If the recent DESI observations DESI:2024mwx ; DESI:2025zgx are correct, there might be a transition from $w<-1$ to $w>-1$ in the DE EoS.
-In this section, we discuss if the (a little bit initial) condition determining if the standard phantom crossing (from $w>-1$ to $w<-1$) or the inverse phantom crossing (from $w<-1$ to $w>-1$) can be realised theoretically in the context of $F(R)$ gravity.
-Note that for future phantom crossing in $F(R)$ gravity, one can consult Bamba:2008hq .
-We show that the (inverse) phantom crossing in the $F(R)$ sector, which is described by the effective pressure and the effective energy density, can occur for almost any $F(R)$ gravity by adjusting the (initial) conditions.
+Let $w$ be the DE equation of state parameter, that is, the ratio of the pressure of the DE fluid with respect to the energy density of the DE fluid.
+If the recent DESI observations DESI:2024mwx ; DESI:2025zgx are correct, there might be a transition from w<1w<-1 to w>1w>-1 in the DE EoS.
+In this section, we discuss if the initial condition determining if the standard phantom crossing (from w>1w>-1 to w<1w<-1) or the inverse phantom crossing (from w<1w<-1 to w>1w>-1) can be realized theoretically in the context of $F(R)$ gravity.
+Therefore, we show that the phantom crossing (or inverse phantom crossing) in the $F(R)$ sector (which is described by the effective pressure and the effective energy density) can occur for almost any $F(R)$ gravity by adjusting the initial conditions.
 The conservation law for a FLRW spacetime is,
 $0=\dot{\rho}+3H\left(\rho+p\right)\,$, (3)
-indicates that when the phantom crossing occurs, since we have $\rho=-p$, then $\dot{\rho}$ must vanish.
+which indicates that when the phantom crossing occurs, since we have $\rho=-p$, then $\dot{\rho}$ must vanish.
 Due to the fact that in the phantom Universe, $\rho<-p$, $\dot{\rho}$ and in the non-phantom Universe, $\rho>p$, there is a crossing from the non-phantom Universe to the phantom one if $\ddot{\rho}>0$ and there is an inverse crossing if $\ddot{\rho}<0$.
@@
 $\rho_{\mathrm{matter}}=\rho_{\mathrm{matter(0)}}a^{-3}=\rho_{\mathrm{matter(0)}}\mathrm{e}^{-3N}\,.$ (8)
-Here $\rho_{\mathrm{matter(0)}}$ is a constant and $a$ is the scale factor, which is given in terms of the $e$-foldings $N$ as $a=\mathrm{e}^{N}$.
-When the phantom crossing in the $f(R)$ sector happens at the cosmic time instance $t=t_{\mathrm{ph}}$, we assume $N$ can be expanded as,
+Here $\rho_{\mathrm{matter(0)}}$ is a constant and $a$ is the scale factor, which is given in terms of the $e$-foldings number $N$ as $a=\mathrm{e}^{N}$.
+When the phantom crossing in the $f(R)$ sector happens at the cosmic time instance $t=t_{\mathrm{ph}}$, we can assume that $N$ can be expanded as follows,
 $N=\sum_{n=0}\frac{H_{n}}{\left(n+1\right)!}\left(t-t_{\mathrm{ph}}\right)^{n+1}\,$, (9)
@@
 $\,+\left(24H_{0}H_{1}+6H_{2}\right)\left(36{H_{0}}^{3}H_{1}-63{H_{0}}^{2}H_{2}-108H_{0}{H_{1}}^{2}-18H_{0}H_{3}-9H_{1}H_{2}\right)$
-×$\,\times f^{\prime\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)$
+$\,\times f^{\prime\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)$
 $\,\left.-324H_{0}\left(4H_{0}H_{1}+H_{2}\right)^{3}f^{\prime\prime\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)\right\}\,.$ (18)
-In the following, we consider the solution of the above equations and their physical meanings. Eq. (14) tells,
+In the following, we consider the solution of the above equations and their physical meanings. Eq. (14) indicates that,
 $H_{1}=-\frac{\kappa^{2}}{2}\rho_{\mathrm{matter(0)}}\,$, (19)
@@
 and noting $\rho_{f(R)}+p_{f(R)}=0$ when the phantom crossing occurs for $f(R)$ sector and $p_{\mathrm{matter}}=0$ for cold matter and baryonic matter.
-By combining (13), (II), and (19), we find the following expression of $H_{2}$,
+By combining (13), (II), and (19), we find the following expression for $H_{2}$,
 $H_{2}=$ $\,-4H_{0}H_{1}-\frac{f\left(12{H_{0}}^{2}+6H_{1}\right)}{36H_{0}f^{\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)}+\frac{\left({H_{0}}^{2}+H_{1}\right)f^{\prime}\left(12{H_{0}}^{2}+6H_{1}\right)}{6H_{0}f^{\prime\prime}\left(12{H_{0}}^{2}+6H_{1}\right)}$
@@
 This could indicate that the phantom crossing or the inverse phantom crossing occurs in a wide class of $F(R)$ gravity models, depending on the initial condition determining $H_{0}$.
-Furthermore, by combining (II) with (II), we can find the expression of $H_{3}$, although we do not write it explicitly because we do not use the expression of $H_{3}$ later.
+Furthermore, by combining (II) with (II), we can find the expression of $H_{3}$ although we do not write it explicitly because we do not use the expression of $H_{3}$ later.
 As an example, we may consider $R^{2}$ gravity,
-$f(R)=\alpha R^{2}\,.$ (23)
-Here $\alpha$ is a constant.
+$f(R)=\alpha R^{2}\,$, (23)
+where $\alpha$ is a constant.
 As is well-known, the $F(R)$ gravity includes a scalar mode.
@@
 $\rho_{{f(R)}\,(2)}=$ $\,\frac{1}{2\kappa^{2}}\left(9{H_{1}}^{2}-\frac{3{H_{0}}^{2}+H_{1}}{6\alpha}\right)\,.$ (24)
-For the standard phantom crossing to occur, the condition is given by
+For the standard phantom crossing to occur, the condition is given by,
 ${H_{0}}^{2}<18\alpha{H_{1}}^{2}-\frac{1}{3}H_{1}=\frac{9\alpha\kappa^{4}}{2}{\rho_{\mathrm{matter(0)}}}^{2}+\frac{\kappa^{2}}{6}\rho_{\mathrm{matter(0)}}\,$, (25)
@@
 III Apparent Phantom Crossing
-There is an idea that instead of considering the transition from $w<-1$ to $w>-1$ of DE, the modification of the dark matter sector may solve the problem in the DESI observations Khoury:2025txd .
-If dark matter density decreases more slowly than $a^{-3}$, as usually predicted from energy conservation, it looks like there could be a phantom crossing in the DE sector because we are considering the total energy density, and we assume the usual $a^{-3}$ behaviour of the dark matter for the analysis of the DESI observations.
+There is an idea that instead of considering the transition from w<1w<-1 to w>1w>-1 of DE, the modification of the dark matter sector may solve the problem in the DESI observations Khoury:2025txd .
+If dark matter density decreases more slowly than $a^{-3}$, as usually predicted from energy conservation, it looks like there could be a phantom crossing in the DE sector because we are considering the total energy density, and we assume the usual $a^{-3}$ behavior of the dark matter for the analysis of the DESI observations.
 If the dark matter is non-relativistic, as in the cold dark matter, the energy density of the dark matter is given by the product $\rho_{\mathrm{DM}}=m_{\mathrm{DM}}n_{\mathrm{DM}}$ of the mass of the dark matter particle, $m_{\mathrm{DM}}$, and the number density of the dark matter particle, $n_{\mathrm{DM}}$.
 If there is neither creation nor annihilation of the dark matter particles, the number density $n_{\mathrm{DM}}$ is proportional to the inverse of the volume in the expanding Universe, $n_{\mathrm{DM}}\propto V^{-1}\propto a^{-3}$.
-If the mass $m_{\mathrm{DM}}$ is time dependent $m_{\mathrm{DM}}(t)$ or $m_{\mathrm{DM}}$ depends on the scalar curvature $R$, the behaviour of the dark matter energy density $\rho_{\mathrm{DM}}$ is different from $a^{-3}$.
-If the mass depends on the scalar curvature, the energy density also includes the derivative terms because the curvature includes the derivative, which makes the situation a little bit complicated.
-In order to make the physical problem at hand well-defined, we assume that the dark matter is a Dirac spinor, whose action is given by
+If the mass $m_{\mathrm{DM}}$ is time dependent $m_{\mathrm{DM}}(t)$ or $m_{\mathrm{DM}}$ depends on the scalar curvature $R$, the behavior of the dark matter energy density $\rho_{\mathrm{DM}}$ is different from $a^{-3}$.
+If the mass depends on the scalar curvature, the energy density also includes the derivative terms because the curvature includes the derivative, which makes the situation somewhat complicated.
+In order to make the physical problem at hand well-defined, we assume that the dark matter is a Dirac spinor, whose action is given by,
 […]
@@
 in the canonical formulation. The energy density is the quantum or statistical average of the Hamiltonian density, as follows,
-[…]
+$\rho_{\mathrm{DM}}=\left<\mathcal{H}\right>=-i\left<\bar{\psi}\gamma^{i}\partial_{i}\psi\right>+m(R)\left<\bar{\psi}\psi\right$>\,. (29)
 The first term corresponds to the kinetic energy and the second term to the rest mass.
 As long as we consider non-relativistic particles, the first term can be neglected by comparing the second term.
-We should also note that $\bar{\psi}\psi\right>$ is nothing but the number density of dark matter $n_{\mathrm{DM}}$.
+We should also note that $\left<\bar{\psi}\psi\right$> is nothing but the number density of dark matter $n_{\mathrm{DM}}$.
 Then we obtain,
@@
 […]
-We assume that $\bar{\psi}\gamma_{i}D_{j}\psi\right>$ and $\left(D_{i}\bar{\psi}\right)\gamma_{j}\psi\right>$ are negligible for the non-relativistic dark matter.
+We assume that $\left<\bar{\psi}\gamma_{i}D_{j}\psi\right$> and $\left<\left(D_{i}\bar{\psi}\right)\gamma_{j}\psi\right$> are negligible for the non-relativistic dark matter.
 Then we find the energy density $\tilde{\rho}_{\mathrm{DM}}$ and the pressure $\tilde{p}_{\mathrm{DM}}$ as follows,
-$\tilde{\rho}_{\mathrm{DM}}=$ $\,m(R)\left<\bar{\psi}\psi\right>+\left\{3\left(\dot{H}+H^{2}\right)+3H\frac{d}{dt}\right\}\left(m^{\prime}(R)\left<\bar{\psi}\psi\right>\right)$
+[…]
 $=$ $\,m(R)n_{\mathrm{DM}}+\left\{3\left(\dot{H}+H^{2}\right)+3H\frac{d}{dt}\right\}\left(m^{\prime}(R)n_{\mathrm{DM}}\right)\,$,
-$\tilde{p}_{\mathrm{DM}}=$$\,\left\{-\left(\dot{H}+3H^{2}\right)+\frac{d^{2}}{dt^{2}}+2H\frac{d}{dt}\right\}\left(m^{\prime}(R)\left<\bar{\psi}\psi\right>\right)$
+[…]
 $=$ $\,\left\{-\left(\dot{H}+3H^{2}\right)+\frac{d^{2}}{dt^{2}}+2H\frac{d}{dt}\right\}\left(m^{\prime}(R)n_{\mathrm{DM}}\right)\,.$ (33)
@@
 $F(R)=R-2\Lambda\left(1-\mathrm{e}^{-\beta\left(\frac{R}{2\Lambda}\right)^{\alpha}}\right)\,$, (35)
-where $\Lambda$ is a constant with dimension $[m]^{2}$ and $\beta$ is a dimensionless constant.
+where $\Lambda$ is a constant with dimension $m]^{2}$ and $\beta$ is a dimensionless constant.
 When $R$ is large, $\beta\left(\frac{R}{2\Lambda}\right)^{\alpha}\gg 1$ if $\alpha>0$, and $\Lambda$ plays the role of the cosmological constant.
@@
 We choose $\frac{\Lambda}{\beta^{\frac{1}{\alpha}}}$ small enough compared with the present value of the scalar curvature.
-This makes $\Lambda$ give the DE in the present Universe.
+This makes $\Lambda$ to realize a DE era in the present Universe.
 In order to also describe the inflation in the early Universe, we may add the term $\gamma R^{2}$ in the $F(R)$ of (35) with a constant $\gamma$.
-Because $f(R)=-2\Lambda\left(1-\mathrm{e}^{-\beta\left(\frac{R}{2\Lambda}\right)^{\alpha}}\right)$, we find
+Because $f(R)=-2\Lambda\left(1-\mathrm{e}^{-\beta\left(\frac{R}{2\Lambda}\right)^{\alpha}}\right)$, we find,
 $f^{\prime}(R)=$ $\,-\alpha\beta\left(\frac{R}{2\Lambda}\right)^{\alpha-1}\mathrm{e}^{-\beta\left(\frac{R}{2\Lambda}\right)^{\alpha}}\,$,
@@
 Then $\rho_{{f(R)}\,(2)}$ in (II) has the following form,
-$\rho_{{f(R)}\,(2)}=$ $\,\frac{3}{\kappa^{2}}\left(-{H_{0}}^{2}H_{1}+{H_{1}}^{2}\right)-\frac{\Lambda^{2}\left(1-\mathrm{e}^{-\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha}}\right)\mathrm{e}^{\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha}}}{3\alpha\beta\kappa^{2}\left\{\left(\alpha-1\right)\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-2}-\alpha\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{2\alpha-2}\right\}}$
+$\rho_{{f(R)}\,(2)}=$ $\,\frac{3}{\kappa^{2}}\left(-{H_{0}}^{2}H_{1}+{H_{1}}^{2}\right)-\frac{2\Lambda\left(1-\mathrm{e}^{-\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha}}\right)\mathrm{e}^{\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha}}}{12\kappa^{2}\frac{\alpha\beta}{2\Lambda}\left\{\left(\alpha-1\right)\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-2}-\alpha\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{2\alpha-2}\right\}}$
+$\,+\frac{\left({H_{0}}^{2}+H_{1}\right)\alpha\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-1}}{2\kappa^{2}\frac{\alpha\beta}{2\Lambda}\left\{\left(\alpha-1\right)\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-2}-\alpha\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{2\alpha-2}\right\}}$
+$\,+\frac{\left(3{H_{0}}^{2}+H_{1}\right)\mathrm{e}^{\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha}}}{6\kappa^{2}\frac{\alpha\beta}{2\Lambda}\left\{\left(\alpha-1\right)\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-2}-\alpha\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{2\alpha-2}\right\}}$
+$=$ $\,\frac{3}{\kappa^{2}}\left(-{H_{0}}^{2}H_{1}+{H_{1}}^{2}\right)-\frac{\Lambda^{2}\left(1-\mathrm{e}^{-\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha}}\right)\mathrm{e}^{\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha}}}{3\alpha\beta\kappa^{2}\left\{\left(\alpha-1\right)\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-2}-\alpha\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{2\alpha-2}\right\}}$
 $\,+\frac{\Lambda\left({H_{0}}^{2}+H_{1}\right)\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-1}}{\kappa^{2}\left\{\left(\alpha-1\right)\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{\alpha-2}-\alpha\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)^{2\alpha-2}\right\}}$
@@
 Then for example, when $\Lambda$ is small enough, $\rho_{{f(R)}\,(2)}\to\frac{3}{\kappa^{2}}\left(-{H_{0}}^{2}H_{1}+{H_{1}}^{2}\right)>0$.
-We should note $H_{1}$ is negative due to (19).
-Then, in the case that $\Lambda$ is small enough, the standard phantom crossing could occur.
-On the other hand, when $\Lambda$ is large enough, we find $\rho_{{f(R)}\,(2)}\to-\frac{\Lambda^{2}}{3\beta^{2}\kappa^{2}}\left(1-\mathrm{e}^{-\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)}\right)<0$, and as long as $-H_{1}<2{H_{0}}^{2}$, the inverse phantom crossing could occur.
+We should note that $H_{1}$ is negative due to (19).
+Then in the case that $\Lambda$ is small enough, the standard phantom crossing could occur.
+On the other hand, when $\Lambda$ is large enough, we find $\rho_{{f(R)}\,(2)}\to-\frac{\Lambda^{2}}{3\beta^{2}\kappa^{2}}\left(1-\mathrm{e}^{-\beta\left(\frac{12{H_{0}}^{2}+6H_{1}}{2\Lambda}\right)}\right)<0$, and as long as $H_{1}<2{H_{0}}^{2}$, the inverse phantom crossing could occur.
 Let us consider more the formulation in Section III. By using (5) and (III), the Friedmann equation in (7) has the following form,
 $0=$ $\,\frac{1}{2\kappa^{2}}\left(R+f(R)\right)-m(R)n_{\mathrm{DM}}$
-+$\,+\left\{3\left(\dot{H}+H^{2}\right)+3H\frac{d}{dt}\right\}\left\{\frac{1}{2\kappa^{2}}\left(1+f^{\prime}(R)\right)-m^{\prime}(R)n_{\mathrm{DM}}\right\}\,.$ (39)
-By using the $e$-foldings $N$ defined by $a=\mathrm{e}^{N}$ by assuming $a=1$ in the present Universe, we find $n_{\mathrm{DM}}=n_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}$, where $n_{\mathrm{DM\,(0)}}$ is the number density of the dark matter particles in the present Universe.
-In terms of $e$-foldings $N$, Eq.
+$\,+\left\{3\left(\dot{H}+H^{2}\right)+3H\frac{d}{dt}\right\}\left\{\frac{1}{2\kappa^{2}}\left(1+f^{\prime}(R)\right)-m^{\prime}(R)n_{\mathrm{DM}}\right\}\,.$ (39)
+By using the $e$-foldings NN defined by $a=\mathrm{e}^{N}$ by assuming $a=1$ in the present Universe, we find $n_{\mathrm{DM}}=n_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}$, where $n_{\mathrm{DM\,(0)}}$ is the number density of the dark matter particles in the present Universe.
+In terms of the $e$-foldings NN, Eq.
 (IV) can be rewritten as,
 $0=$ $\,\frac{1}{2\kappa^{2}}\left(R+f(R)\right)-m(R)n_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}$
-+$\,+\left\{3\left(HH^{\prime}+H^{2}\right)+3H^{2}\frac{d}{dN}\right\}\left\{\frac{1}{2\kappa^{2}}\left(1+f^{\prime}(R)\right)-m^{\prime}(R)n_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}\right\}\,,$ (40)
+$\,+\left\{3\left(HH^{\prime}+H^{2}\right)+3H^{2}\frac{d}{dN}\right\}\left\{\frac{1}{2\kappa^{2}}\left(1+f^{\prime}(R)\right)-m^{\prime}(R)n_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}\right\}\,$, (40)
 where $H^{\prime}\equiv\frac{dH}{dN}$ and we used the relation $\frac{d}{dt}=H\frac{d}{dN}$.
@@
 $=$ $\,-\left\{3\left(HH^{\prime}+H^{2}\right)\Phi(N)+3H^{2}\frac{d\Phi(N)}{dN}\right\}\,$, (41)
-Because,
+Since,
 $\Psi^{\prime}=\left(24HH^{\prime}+6{H^{\prime}}^{2}+6HH^{\prime\prime}\right)\Phi+3m(R)n_{\mathrm{DM\,(0)}}\mathrm{e}^{-3N}\,$, (42)
@@
 $m\left(R\left(N\right)\right)=\frac{\mathrm{e}^{3N}}{3n_{\mathrm{DM\,(0)}}}\left(\Psi^{\prime}-\left(24HH^{\prime}+6{H^{\prime}}^{2}+6HH^{\prime\prime}\right)\Phi\right)\,$, (43)
-Using the expression of $N=N(R)$, we find $m(R)$ as a function of $R$. Then by the definition of $\Psi$, we find,
+Using the expression of $N=N(R)$, we find $m(R)$ as a function of $R$. Then by using the definition of $\Psi$, we find,
 $f(R)=-R+2\kappa^{2}\Psi\left(N\left(R\right)\right)+2\kappa^{2}m\left(R\right)\mathrm{e}^{-3N(R)}\,.$ (44)
-Then we find $f(R)$ as a function of $R$. First, we consider a very trivial example,
+Then we can find $f(R)$ as a function of $R$. First, we consider a very trivial example,
 $H=H_{0}\,,\quad\Phi=\Phi_{0}\,$, (45)
-with constants $H_{0}$ and $\Phi_{0}$ to check the consistency in the above formulation. The spacetime is, of course, the de Sitter spacetime. Then Eqs. (IV), (43), and (44) tell,
+with constants $H_{0}$ and $\Phi_{0}$ to check the consistency in the above formulation. The spacetime is, of course, the de Sitter spacetime. Then Eqs. (IV), (43), and (44) indicate that,
 $\Psi=-3{H_{0}}^{2}\Phi_{0}\,,\quad m(R)=0\,,\quad f(R)=-12{H_{0}}^{2}-6\kappa^{2}\Psi_{0}\,.$ (46)
-Then we obtained the model with a cosmological constant, where $f(R)$ is a constant and there is no matter $m(R)=0$, as expected.
-Anyway, we find the above formulation using Eqs.
+Hence we obtained the model with a cosmological constant, where $f(R)$ is a constant and there is no matter $m(R)=0$, as expected.
+We can find the above formulation using Eqs.
 (IV), (43), and (44) may work.
-Second, as a less non-trivial example, we consider the case that
+Second, as a less non-trivial example, we consider the case that,
 $H^{2}={H_{0}}^{2}+{H_{1}}^{2}\mathrm{e}^{-3N}\,$, (47)
@@
 $f(R)=$ $\,\left(-1+\frac{2\kappa^{2}\Phi_{0}}{n_{\mathrm{DM\,(0)}}}\right)R+6\kappa^{2}{H_{0}}^{2}\Phi_{0}\left(-1+\frac{4}{n_{\mathrm{DM\,(0)}}}\right)\,.$ (48)
-Therefore, we obtain Einstein’s gravity with a cosmological constant consistently, although the gravitational coupling $\kappa$ is effectively modified as $\kappa^{2}\to\frac{n_{\mathrm{DM\,(0)}}}{\Phi_{0}}$ because $F(R)=R+f(R)=\frac{2\kappa^{2}\Phi_{0}}{n_{\mathrm{DM\,(0)}}}R+6\kappa^{2}{H_{0}}^{2}\Phi_{0}\left(-1+\frac{4}{n_{\mathrm{DM\,(0)}}}\right)$.
+Therefore, we obtain Einstein’s gravity with a cosmological constant consistently although the gravitational coupling $\kappa$ is effectively modified as $\kappa^{2}\to\frac{n_{\mathrm{DM\,(0)}}}{\Phi_{0}}$ because $F(R)=R+f(R)=\frac{2\kappa^{2}\Phi_{0}}{n_{\mathrm{DM\,(0)}}}R+6\kappa^{2}{H_{0}}^{2}\Phi_{0}\left(-1+\frac{4}{n_{\mathrm{DM\,(0)}}}\right)$.
 We now consider a more realistic and non-trivial model.
@@
 $\dot{\rho}^{\mathrm{eff}}_{\mathrm{DE}}=0\,,\quad\ddot{\rho}^{\mathrm{eff}}_{\mathrm{DE}}<0\,.$ (49)
-The $e$-foldings $N$ is a monotonically increasing function of the cosmic time $t$, and the above condition (49) can be rewritten in terms of $N$, as follows,
+The $e$-foldings NN is a monotonically increasing function of the cosmic time $t$, and the above condition (49) can be rewritten in terms of $N$, as follows,
 ${\rho^{\mathrm{eff}}_{\mathrm{DE}}}^{\prime}=0\,,\quad{\rho^{\mathrm{eff}}_{\mathrm{DE}}}^{\prime\prime}<0\,.$ (50)
@@
 $\rho^{\mathrm{eff}}_{\mathrm{DE}}=\Lambda_{0}+\frac{\Lambda_{1}}{\left(\mathrm{e}^{-3N}-\mathrm{e}^{-3N_{0}}\right)^{2}+C^{2}}\,.$ (51)
-Here $\Lambda_{0}$ and $\Lambda_{1}$ are constants with the dimension of $[m]^{4}$, and $C$ is a dimensionless constant.
+Here $\Lambda_{0}$ and $\Lambda_{1}$ are constants with the dimension of $m]^{4}$, and $C$ is a dimensionless constant.
 The effective energy density $\rho^{\mathrm{eff}}_{\mathrm{DE}}$ has a peak at $N=N_{0}$, which corresponds to the $e$-foldings that the inverse phantom crossing occurs.
@@
 As in Elizalde:2011ds ; Elizalde:2010ts , we consider the model where the Hubble rate $H$ oscillates and therefore the Universe iterates the phantom crossing and the inverse phantom crossing.
-In this section, we consider such an oscillation for the total energy density or the behaviour of $H$.
+In this section, we consider such a oscillation for the total energy density or the behavior of $H$.
 A simple model of this sort is given by,
 $H^{2}={H_{0}}^{2}\left(1+C\sin\left(\omega N\right)\right)\,$, (58)
-where $H_{0}$, $C$, and $\omega$ are constants, and we assume $10<C<$ 1 and $\omega>0$. Then we find,
+where $H_{0}$, $C$, and $\omega$ are constants, and we assume $0<C<1$ and $\omega>0$. Then we find,
 $HH^{\prime}=\frac{{H_{0}}^{2}C\omega}{2}\cos\left(\omega N\right)\,.$ (59)
-Because we are considering the expanding Universe, where $H>0$, when $-\frac{\pi}{2}+2n\pi<\omega N<\frac{\pi}{2}+2n\pi$ with an integer $n$, $H^{\prime}$ is positive and therefore, the Universe is in the phantom phase.
+Because we are considering the expanding Universe, where $H>0$, when $\frac{\pi}{2}+2n\pi<\omega N<\frac{\pi}{2}+2n\pi$ with an integer $n$, $H^{\prime}$ is positive and therefore, the Universe is in the phantom phase.
 On the other hand, when $\frac{\pi}{2}+2n\pi<\omega N<\frac{3\pi}{2}+2n\pi$ with an integer $n$, $H^{\prime}$ is negative and therefore, the Universe is in the non-phantom phase.
@@
 Under the assumption (58), the scalar curvature $R$ is given by
-$R=12H^{2}+6HH^{\prime}=3{H_{0}}^{2}\left(4+\sqrt{16+\omega^{2}}C\sin\left(\omega N+\theta_{0}\right)\right)\,.$ (60)
-Here $\theta_{0}$ is defined by $\tan\theta_{0}=\frac{\omega}{4}$. Eq. (60) can be solved with respect to $N$,
+$R=$  $\,12H^{2}+6HH^{\prime}=H_{0}^{2}\left(12+12C\sin\left(\omega N\right)+3C\omega\cos\left(\omega N\right)\right)$
+$=$  $\,3{H_{0}}^{2}\left\{4+\sqrt{16+\omega^{2}}C\left(\frac{4}{\sqrt{16+\omega^{2}}}\sin\left(\omega N\right)+\frac{\omega}{\sqrt{16+\omega^{2}}}\cos\left(\omega N\right)\right)\right\}$
+$=$  $\,3{H_{0}}^{2}\left(4+\sqrt{16+\omega^{2}}C\sin\left(\omega N+\theta_{0}\right)\right)\,.$ (60)
+Here $\theta_{0}$ is defined by $\tan\theta_{0}=\frac{\omega}{4}$. Eq. (V) can be solved with respect to $N$,
 $N=-\frac{\theta_{0}}{\omega}\sin^{-1}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\,.$ (61)
-Here $sin^{-1}$ is the inverse function of $\sin$ function. The $e$-foldings $N$ is explicitly given as a function of $R$. Furthermore, we find,
-$\sin\left(\omega N\right)=$ $\,\frac{4}{\sqrt{16+\omega^{2}}}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)-\frac{\omega}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}\,,$
-$\cos\left(\omega N\right)=$ $\,\frac{4}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}+\frac{\omega}{\sqrt{16+\omega^{2}}}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\,,$ (62)
+Here $sin^{-1}$ is the inverse function of $\sin$ function. The $e$-foldings NN is explicitly given as a function of $R$. Furthermore, we find,
+$\sin\left(\omega N\right)=$ $\,\sin\left(\omega N+\theta_{0}-\theta_{0}\right)=\sin\left(\omega N+\theta_{0}\right)\cos\left(\theta_{0}\right)-\cos\left(\omega N+\theta_{0}\right)\sin\left(\theta_{0}\right)$
+$=$ $\,\frac{4}{\sqrt{16+\omega^{2}}}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)-\frac{\omega}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}\,$,
+$\cos\left(\omega N\right)=$ $\,\cos\left(\omega N+\theta_{0}-\theta_{0}\right)=\cos\left(\omega N+\theta_{0}\right)\cos\left(\theta_{0}\right)+\sin\left(\omega N+\theta_{0}\right)\sin\left(\theta_{0}\right)$
+$=$ $\,\frac{4}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}+\frac{\omega}{\sqrt{16+\omega^{2}}}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\,$, (62)
 which we will use shortly. Again, we use Eqs. (IV), (43), and (44). Then by assuming $\Phi=\Phi_{0}$, we find,
 $\Psi(N)=$ $\,-3\Phi_{0}{H_{0}}^{2}\left(1+C\sin\left(\omega N\right)+\frac{C\omega}{2}\cos\left(\omega N\right)\right)\,$,
-$m(N)=$ $\,-\frac{\Phi_{0}{H_{0}}^{2}C\mathrm{e}^{3N}}{n_{\mathrm{DM\,(0)}}}\left\{\frac{9\omega^{2}}{2}\sin\left(\omega N\right)-15\omega\cos\left(\omega N\right)\right\}\,,$
-$f=$ $\,-R-2\kappa^{2}\Phi_{0}{H_{0}}^{2}\left[3+C\left\{\left(3+\frac{9\omega^{2}}{2n_{\mathrm{DM\,(0)}}}\right)\sin\left(\omega N\right)\right.\right.$
+$m(N)=$ $\,\frac{\Phi_{0}{H_{0}}^{2}\mathrm{e}^{3N}}{3n_{\mathrm{DM\,(0)}}}\left\{-3\left(C\omega\cos\left(\omega N\right)-\frac{C\omega^{2}}{2}\sin\left(\omega N\right)\right)-12C\omega\cos\left(\omega N\right)+3C\omega^{2}\sin\left(\omega N\right)\right\}$
+$=$ $\,-\frac{\Phi_{0}{H_{0}}^{2}C\mathrm{e}^{3N}}{n_{\mathrm{DM\,(0)}}}\left\{\frac{9\omega^{2}}{2}\sin\left(\omega N\right)-15\omega\cos\left(\omega N\right)\right\}\,$,
+$f=$ $\,-R+2\kappa^{2}\left\{-3\Phi_{0}{H_{0}}^{2}\left(1+C\sin\left(\omega N\right)+\frac{C\omega}{2}\cos\left(\omega N\right)\right)\right\}$
+$\,-\frac{2\kappa^{2}\Phi_{0}{H_{0}}^{2}C}{n_{\mathrm{DM\,(0)}}}\left\{\frac{9\omega^{2}}{2}\sin\left(\omega N\right)-15\omega\cos\left(\omega N\right)\right\}$
+[…]
 $\,\left.\left.+\left(\frac{3}{2}-\frac{15}{n_{\mathrm{DM\,(0)}}}\right)\omega\cos\left(\omega N\right)\right\}\right]\,.$ (63)
 Then by using (61), we find the $R$ dependence of $m=m(R)$ and $f=f(R)$,
-$m(R)=$ $\,-\frac{\Phi_{0}{H_{0}}^{2}C\mathrm{e}^{-\frac{3\theta_{0}}{\omega}\sin^{-1}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)}}{n_{\mathrm{DM\,(0)}}\sqrt{16+\omega^{2}}}\left\{3\omega^{2}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\right.$
+[…]
+$\,\left.-\frac{\omega}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}\right\}-15\omega\left\{\frac{4}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}\right$.
+$\,\left.\left.+\frac{\omega}{\sqrt{16+\omega^{2}}}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\right\}\right]\,$,
+$=$ $\,-\frac{\Phi_{0}{H_{0}}^{2}C\mathrm{e}^{-\frac{3\theta_{0}}{\omega}\sin^{-1}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)}}{n_{\mathrm{DM\,(0)}}\sqrt{16+\omega^{2}}}\left\{3\omega^{2}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\right$.
 $\,\left.-\left(\frac{9\omega^{3}}{2}+60\omega\right)\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}\right\}\,$,
-$f(R)=$ $\,-R-2\kappa^{2}\Phi_{0}{H_{0}}^{2}\left[3+\frac{C}{\sqrt{16+\omega^{2}}}\left\{\left(12+\frac{3}{2}\omega^{2}+\frac{3\omega^{2}}{n_{\mathrm{DM\,(0)}}}\right)\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\right.\right.$
+[…]
+$\,\left.-\frac{\omega}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}\right\}$
+$\,+\left(\frac{3}{2}-\frac{15}{n_{\mathrm{DM\,(0)}}}\right)\omega\left\{\frac{4}{\sqrt{16+\omega^{2}}}\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}\right$.
+$\,\left.\left.+\frac{\omega}{\sqrt{16+\omega^{2}}}\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)\right\}\right]$
+[…]
 $\,\left.\left.+\left(3\omega-\frac{120\omega+9\omega^{3}}{2n_{\mathrm{DM\,(0)}}}\right)\sqrt{1-\left(\frac{R-12{H_{0}}^{2}}{3{H_{0}}^{2}C\sqrt{16+\omega^{2}}}\right)^{2}}\right\}\right]$
@@
 Then the particles corresponding to the scalar mode might become dark matter, which may explain the DESI observation.
-The action of the $F(R)$ gravity is given by replacing the scalar curvature $R$ in the Einstein-Hilbert action, which is,
+The action of the $F(R)$ gravity is given by replacing the scalar curvature $R$ in the Einstein-Hilbert action which is,
 $S_{\mathrm{EH}}=\int d^{4}x\sqrt{-g}\left(\frac{R}{2\kappa^{2}}+\mathcal{L}_{\mathrm{matter}}\right)\,$, (65)
@@
 We obtain $A=R$ by the variation of the action with respect to $A$ and by substituting the obtained equation $A=R$ into the action (70), we find that the action in (66) is reproduced.
-If we rescale the metric by a kind of scale transformation,
+If we rescale the metric by a kind of a scale transformation,
 $g_{\mu\nu}\to\mathrm{e}^{\sigma}g_{\mu\nu}\,,\quad\sigma=-\ln F^{\prime}(A)\,$, (71)
@@
 Therefore, we may estimate $m_{\sigma}$ as $m_{\sigma}\sim 10^{13}\,\mathrm{GeV}$, which is large enough.
-Anyway, the model (74) simply describes both the inflation and the DE.
+In any case, the model (74) simply describes both the inflation and the DE.
 VII Towards a Realistic Model: A Theoretical Approach
@@
 Then, in the thermal equilibrium, the scalaron could be created by the interaction.
-Because the coupling is given by the gravitational coupling, which is the inverse of the Planck mass, the interaction is very weak, and therefore, a sufficient amount of the scalaron particles could be created, and the mass of the scalaron must be small enough.
+Because the coupling is given by the gravitational coupling, which is the inverse of the Planck mass, the interaction is very weak and therefore a sufficient amount of the scalaron particles could be created, and the mass of the scalaron must be small enough.
 Eq.
@@
 There might also be gravitational particle production.
-We should note that the DESI observation corresponds to the redshift region $3.50<z<$ 3.5.
-If we choose the parameter $\alpha$ in the model in (74) to be $\alpha\sim\left(10^{13}\,\mathrm{GeV}\right)^{-2}$ so that the model could also describe the inflation, Eq.
+We should note that the DESI observation corresponds to the redshift region $0<z<3.5$.
+If we choose the parameter $\alpha$ in the model in (74) to be $\alpha\sim\left(10^{13}\,\mathrm{GeV}\right)^{-2}$ so that the model could also describe the inflationary era, Eq.
 (VI) tells that in the DESI redshift, the mass is already heavy enough, ${m_{\sigma}}^{2}\sim\frac{3}{4\alpha}$, and therefore the inverse phantom crossing could not occur.
 In order to avoid this problem, we may choose $1/\alpha$ to be small enough.
-In the laboratory, the energy density of the redshift region $3.50<z<$ 3.5 can be easily realised, and the fifth force experiment excludes such a light mode.
+In the laboratory, the energy density of the redshift region $0<z<3.5$ can be easily realized, and the fifth force experiment excludes such a light mode.
 The coupling between the scalaron and the matter shifts the mass due to the Chameleon mechanism, but the scalaron could be light enough to be excluded from the experiments.
-In order to avoid the problem, we may modify the model so that $\alpha$ could depend on the scalar curvature.
-The problem might be solved if $1/\alpha$ is large enough, even in the redshift region $3.50<z<$ 3.5 to be consistent with the fifth force experiment, and after that $1/\alpha$ further increases to generate the apparent phantom crossing.
+In order to avoid this problem, we may modify the model so that $\alpha$ could depend on the scalar curvature.
+The problem might be solved if $1/\alpha$ is large enough even in the redshift region $0<z<3.5$ to be consistent with the fifth force experiment and after that $1/\alpha$ further increases to generate the apparent phantom crossing.
 VIII Some Explicit Examples of Oscillating DE with Phantom Crossing
-Now let us consider some realistic models of $F(R)$ gravity DE which we will extensively study using numerical methods, and we will obtain the cosmological parameters of these models.
-We will be mainly interested in the DE density and DE EoS parameter, which are constrained by the data, and we will numerically show that the models exhibit late-time EoS oscillations and a phantom crossing.
+Now let us consider some realistic models of $F(R)$ gravity DE which we will extensively study using numerical methods and we will obtain the cosmological parameters of these models.
+We will be mainly interested in the DE density and DE EoS parameter which are constrained by the data and we will numerically show that the models exhibit late-time EoS oscillations and a phantom crossing.
 Let us consider the $F(R)$ gravity action in the presence of perfect matter fluids in (66).
@@
 where $\rho_{\mathrm{total}}$ stands for the total energy density of the total effective cosmological fluid, which includes the $F(R)$ gravity contribution, and $p_{\mathrm{total}}$ denotes the total pressure.
-The total cosmological fluid is composed of three parts: the cold dark matter one ($\rho_{\mathrm{matter}}$), the radiation ($\rho_{\mathrm{r}}$) and the geometric part of $F(R)$ gravity ($\rho_{\mathrm{DE}}$).
+The total cosmological fluid is composed by three parts, the cold dark matter one ($\rho_{\mathrm{matter}}$), the radiation ($\rho_{\mathrm{r}}$) and the geometric part of $F(R)$ gravity ($\rho_{\mathrm{DE}}$).
 Therefore, we have, $\rho_{\mathrm{total}}=\rho_{\mathrm{matter}}+\rho_{\mathrm{r}}+\rho_{\mathrm{DE}}$ and in addition $p_{\mathrm{total}}=p_{\mathrm{matter}}+p_{\mathrm{r}}+p_{\mathrm{DE}}$.
@@
 $1+z=\frac{1}{a}\,$, (83)
-and in addition, we introduce the statefinder parameter $y_{H}(z)$ Hu:2007nk ; Bamba:2012qi ; Nojiri:2017ncd ,
+and in addition we introduce the statefinder parameter $y_{H}(z)$ Hu:2007nk ; Bamba:2012qi ; reviews1 ,
 $y_{H}(z)=\frac{\rho_{\mathrm{DE}}}{\rho_{\mathrm{matter}}^{(0)}}=\frac{H^{2}}{{m_{s}}^{2}}-(1+z)^{3}-\chi(1+z)^{4},$ (84)
@@
 where the dimensionless functions $J_{1}$ , $J_{2}$ , $J_{3}$ are defined as,
-$J_{1}=$ $\,\frac{1}{(z+1)}\left(-3-\frac{1}{y_{H}+(z+1)^{3}+\chi(z+1)^{4}}\frac{1-F_{R}}{6{m_{s}}^{2}F_{RR}}\right)\,,$ (86)
-$J_{2}=$ $\,\frac{1}{(z+1)^{2}}\left(\frac{1}{y_{H}+(z+1)^{3}+\chi(z+1)^{4}}\frac{2-F_{R}}{3{m_{s}}^{2}F_{RR}}\right)\,,$ (87)
-$J_{3}=$ $\,-3(z+1)-\frac{(1-F_{R})\left((z+1)^{3}+2\chi(z+1)^{4}\right)+(R-F)/\left(3{m_{s}}^{2}\right)}{(z+1)^{2}\left(y_{H}+(z+1)^{3}+\chi(z+1)^{4}\right)}\frac{1}{6{m_{s}}^{2}F_{RR}}\,,$ (88)
+$J_{1}=\frac{1}{(z+1)}\Big(-3-\frac{1}{y_{H}+(z+1)^{3}+\chi(z+1)^{4}}\frac{1-F_{R}}{6{m_{s}}^{2}F_{RR}}\Big)\,$, (86)
+$J_{2}=\frac{1}{(z+1)^{2}}\Big(\frac{1}{y_{H}+(z+1)^{3}+\chi(z+1)^{4}}\frac{2-F_{R}}{3{m_{s}}^{2}F_{RR}}\Big)\,$, (87)
+$J_{3}=-3(z+1)-\frac{(1-F_{R})((z+1)^{3}+2\chi(z+1)^{4})+(R-F)/(3{m_{s}}^{2})}{(z+1)^{2}(y_{H}+(z+1)^{3}+\chi(z+1)^{4})}\frac{1}{6{m_{s}}^{2}F_{RR}}\,$, (88)
 and furthermore $F_{RR}=\frac{\partial^{2}F}{\partial R^{2}}$. Moreover, the Ricci scalar is,
 $R=12H^{2}-6HH_{z}(1+z)\,$, (89)
-and in terms of $y_{H}$, we obtain,
+and in terms of $y_{H}$ we obtain,
 $R(z)=3\,{m_{s}}^{2}\left(-(z+1)\,\frac{dy_{H}(z)}{dz}+4y_{H}(z)+(1+z)^{3}\right)\,.$ (90)
 We will solve Eq. (85) numerically for the late-time redshift interval $z=[0,10]$, using the following initial conditions at the final redshift $z_{f}=10$ Bamba:2012qi ,
-$y_{H}(z_{f})=\frac{\Lambda}{3{m_{s}}^{2}}\left(1+\frac{1+z_{f}}{1000}\right)\,,\quad\left.\frac{dy_{H}(z)}{dz}\right|_{z=z_{f}}=\frac{1}{1000}\frac{\Lambda}{3{m_{s}}^{2}}\,,$ (91)
+$y_{H}(z_{f})=\frac{\Lambda}{3{m_{s}}^{2}}\Big(1+\frac{1+z_{f}}{1000}\Big)\ ,\ \frac{dy_{H}(z)}{dz}\Big|_{z=z_{f}}=\frac{1}{1000}\frac{\Lambda}{3{m_{s}}^{2}},$ (91)
 where $\Lambda\simeq 11.895\times 10^{-67}\,\mathrm{eV}^{2}$. We can express the physical cosmological quantities using the statefinder $y_{H}(z)$, and we have,
@@
 and the Ricci scalar is,
-$R(z)=3{m_{s}}^{2}\left(4y_{H}(z)-(z+1)\frac{dy_{H}(z)}{dz}+(z+1)^{3}\right)\,,$ (93)
+$R(z)=3{m_{s}}^{2}\Big(4y_{H}(z)-(z+1)\frac{dy_{H}(z)}{dz}+(z+1)^{3}\Big),$ (93)
 while the DE density parameter $\Omega_{\mathrm{DE}}(z)$ is,
@@
 $q(z)=-1-\frac{\dot{H}}{H^{2}}=-1-(z+1)\frac{H^{\prime}(z)}{H(z)},$ (97)
-where the “prime” denotes differentiation with respect to the redshift. In addition, and for comparison reasons, let us quote here the Hubble rate for the $\Lambda$CDM model,
+[…]
 $H_{\Lambda}(z)=H_{0}\sqrt{\Omega_{\Lambda}+\Omega_{\mathrm{matter}}(z+1)^{3}+\Omega_{\mathrm{r}}(z+1)^{4}},$ (98)
 where $\Omega_{\Lambda}\simeq 0.68136$ and $\Omega_{\mathrm{matter}}\simeq 0.3153$ and $H_{0}\simeq 1.37187\times 10^{-33}\,\mathrm{eV}$ according to the latest 2018 Planck data Planck:2018vyg .
-Now let us focus on the phenomenology of two interesting models which exhibit remarkable behaviour at late times.
+Now let us focus on the phenomenology of two interesting models which exhibit remarkable behavior at late times.
 We start off with the following model,
 $F(R)=R+\frac{R^{2}}{M^{2}}-\frac{b}{c+\exp(-R/R_{0})},$ (99)
-with $b$ and $R_{0}$ are free parameters with mass dimensions $[m]^{2}$ and also $c$ is a dimensionless parameter.
-We obtained a viable phenomenology for the choice $b=20\Lambda,$ $c=2$, and $R_{0}={m_{s}}^{2}/0.00091$.
-In Fig.
-1 we plot the deceleration parameter for this model (red curve) compared with the $\Lambda$CDM (black curve), and the DE EoS for various redshift ranges.
-As can be seen, the DE EoS is strongly oscillating at late times, and the model ends up in a quintessential regime.
-The DE energy density parameter and the DE EoS parameter at $z=0$ are for this model $\Omega_{\mathrm{DE}}=0.6918$ and $w_{\mathrm{DE}}=-0.9974$ which comply with the latest Planck constraints.
+[…]
 Figure 1: Plots of the deceleration parameter $q(z)$ (upper left plot) and the DE EoS parameter $w_{\mathrm{DE}}(z)$ (right and bottom plot) as functions of the redshift for the model Eq.
@@
 $F(R)=R+\frac{R^{2}}{M^{2}}-\alpha\frac{b(R/R_{0})^{n}}{c(R/R_{0})^{n}+d},$ (100)
-with $b$, $c$, $d$, $n$ being dimensionless parameters and $\alpha$, $R_{0}$ are also free parameters with mass dimensions $[m]^{2}$.
-With the choice $\alpha=1.4\Lambda$, $b=1$, $c=0.2$, $d=0.04$, $R_{0}={m_{s}}^{2}$, and $n=0.3$, we obtain an interesting DE oscillating behaviour which is also phenomenologically viable.
-Specifically, in Fig.
-2, we plot the deceleration parameter (red curve) compared with the $\Lambda$CDM (black curve), and also the DE EoS for various redshift ranges, focusing on small redshifts.
-As can be seen in this case, too, the DE EoS is strongly oscillating at late times and at small redshifts before the present day, and the model ends up in a quintessential regime in this case, too.
-The DE energy density parameter and the DE EoS parameter at $z=0$ for this model are $\Omega_{\mathrm{DE}}=0.6851$ and $w_{\mathrm{DE}}=-0.9887$ which are again compatible with the latest Planck constraints.
+[…]
 Figure 2: Plots of the deceleration parameter $q(z)$ (upper left plot) and the DE EoS parameter $w_{\mathrm{DE}}(z)$ (right and bottom plot) as functions of the redshift for the model of Eq.
 (100) with $\alpha=1.4\Lambda$, $b=1$, $c=0.2$ , $d=0.04$, $R_{0}={m_{s}}^{2}$, and $n=0.3$.
-Thus in this section, using a numerical approach, we demonstrated that it is possible to obtain oscillating DE eras using rather simple $F(R)$ gravity models.
+Thus in this section using a numerical approach, we demonstrated that it is possible to obtain oscillating DE eras using rather simple $F(R)$ gravity models.
 Note that the models presented in this section were also extensively examined regarding their phenomenology in Ref.
@@
 IX Conclusions
-In this paper, we studied the possibility of having dynamical DE which crosses the phantom divide line and oscillates at late times.
+In this paper we studied the possibility of having dynamical DE which crosses the phantom divide line and oscillates at late times.
 We examined this problem both theoretically and numerically and proved that $F(R)$ gravity provides a consistent phenomenological framework for naturally describing the DE era without resorting to phantom fields.
@@
 We also discussed the apparent phantom crossing issue, further extending the considerations made.
-Furthermore, we developed the theoretical study of several models that may exhibit phantom crossing and discussed from a theoretical viewpoint these models and their DE behaviour, in a qualitative way, though.
-Moreover, several oscillating forms of the Hubble rate were considered, in terms of the $e$-foldings number and their realisation in $F(R)$ gravity was also considered.
-In addition, we considered the effects of a scalaron as dark matter with a curvature-dependent mass on the DE era.
-We also studied in a quantitative way the predictions on the DE era of two explicit $F(R)$ gravity models, one of which is also compatible with the DESI data, apart from the Planck data.
-We studied the behaviour of the deceleration parameter and the DE EoS, and showed that it is possible to obtain a phantom crossing EoS and, in addition, an oscillating DE EoS at late times.
-Thus, we proved that indeed $F(R)$ gravity offers a consistent framework which can generate a phantom to quintessence transition in a natural way without phantom fields, and moreover, some of the models realise DE oscillations, again built in the model, not artificially added.
+Furthermore, we developed the theoretical study of several models that may exhibit phantom crossing and discussed from a theoretical viewpoint these models and their DE behavior, in a qualitative way though.
+Moreover, several oscillating forms of the Hubble rate were considered, in terms of the $e$-foldings number and their realization in $F(R)$ gravity was also considered.
+In addition, we studied the effects of a scalaron as dark matter with a curvature-dependent mass on the DE era.
+We also studied in a quantitative way the predictions on the DE era of two explicit $F(R)$ gravity models, one of which is also compatible with the DESI data apart from the Planck data.
+We considered the behavior of the deceleration parameter and the DE EoS and showed that it is possible to obtain a phantom crossing EoS and in addition an oscillating DE EoS at late times.
+Thus we proved that indeed $F(R)$ gravity offers a consistent framework which can generate a phantom to quintessence transition in a natural way without phantom fields and moreover some of the models realize DE oscillations, again built in the model, not artificially added.
 Now the question is why should $F(R)$ gravity should be considered a viable extension of Einstein’s gravity.
-The answer is simple: Einstein’s gravitational Lagrangian density is described simply by $R$, so an Occam’s razor approach is that an $F(R)$ gravity extends it.
+The answer is simple, Einstein’s gravitational Lagrangian density is described simply by $R$, so an Occam’s razor approach is that an $F(R)$ gravity extends it.
 Of course, someone may also seek if such oscillating and phantom divide crossing phenomena occur in other gravities, like Einstein-Gauss-Bonnet theories, but the natural extension of Einstein’s gravity is $F(R)$ gravity and it enjoys an elevated role among all modified gravity models.
 Acknowledgements
-We are indebted to Taishi Katsuragawa for discussions. This work was partially supported by the program Unidad de Excelencia Maria de Maeztu CEX2020-001058-M, Spain (S.D.O).
+This work was partially supported by the program Unidad de Excelencia Maria de Maeztu CEX2020-001058-M, Spain (S.D.O).